### v1

# TempoNet: Slack-Quantized Transformer-Guided Reinforcement Scheduler for Adaptive Deadline-Centric Real-Time Dispatchs

Abstract
Real-time schedulers must reason about tight deadlines under strict compute budgets. We present TempoNet, a reinforcement learning scheduler that pairs a permutation-invariant Transformer with a deep Q-approximation. An Urgency Tokenizer discretizes temporal slack into learnable embeddings, stabilizing value learning and capturing deadline proximity. A latency-aware sparse attention stack with blockwise top-$k$ selection and locality-sensitive chunking enables global reasoning over unordered task sets with near-linear scaling and sub-millisecond inference. A multicore mapping layer converts contextualized Q-scores into processor assignments through masked-greedy selection or differentiable matching. Extensive evaluations on industrial mixed-criticality traces and large multiprocessor settings show consistent gains in deadline fulfillment over analytic schedulers and neural baselines, together with improved optimization stability. Diagnostics include sensitivity analyses for slack quantization, attention-driven policy interpretation, hardware-in-the-loop and kernel micro-benchmarks, and robustness under stress with simple runtime mitigations; we also report sample-efficiency benefits from behavioral-cloning pretraining and compatibility with an actor-critic variant without altering the inference pipeline. These results establish a practical framework for Transformer-based decision making in high-throughput real-time scheduling.
Keywords Reinforcement Learning, Real-Time Systems, Transformer Models, Attention Mechanisms, Scheduling under Uncertainty, Resource Allocation, Embedded AI
1 Introduction
Real-time systems require schedulers that make correct, low-latency decisions under dynamic workloads. Classical policies such as Rate Monotonic and Earliest Deadline First provide strong guarantees under ideal assumptions but degrade under bursty loads or uncertain execution times, motivating multicore strategies and empirical studies beyond idealized conditions [33, 2].
Data-driven approaches address these limitations by learning policies from interaction data. RL has shown promise for cloud orchestration, job-shop scheduling, and cluster placement [40, 9, 47, 22, 24], with offline RL and imitation learning improving sample efficiency in constrained domains [35].
However, many RL schedulers rely on sequence encodings or fixed-size vectors, introducing order dependence and limiting generalization [19, 38]. Set- and graph-based models mitigate these issues, yet integrating them under strict sub-millisecond inference budgets remains challenging [23]. Transformers enable global reasoning via attention, making them attractive for scheduling where cross-task interactions matter. Multi-head attention supports parallel pairwise modeling [39, 4, 8]. Sequence-modeling approaches such as Decision Transformer excel in offline RL but depend on ordered histories and causal masking, unsuitable for unordered sets and tight latency constraints [5]. Online attention-based agents improve representation capacity [17], while hardware co-designs enhance throughput without guaranteeing tail-latency bounds [32]. Sparse and selective attention methods, including explicit key selection and block-sparse routing, offer compression strategies, and RL-guided quantization reduces runtime cost; however, adapting these techniques to hard real-time value-based schedulers requires careful co-design of representation, sparsification, and mapping strategies [48, 30, 21, 36, 51] We introduce TempoNet, a value-based RL scheduler designed for predictable, low-latency operation with global reasoning. TempoNet combines three design choices: a slack-quantized token representation that discretizes continuous slack into learnable embeddings, reducing gradient variance and focusing attention on deadline-aware groups; a compact, permutation-invariant attention encoder with shallow depth, narrow width, and sparsification via block Top-$k$ and locality-aware chunking for near-linear scaling; and multicore mapping layers that translate per-token Q-values into core assignments under latency and migration constraints using masked-greedy or bipartite matching variants. TempoNet is trained with stable value-based updates and engineered exploration schedules for robustness under overload. Experiments on uniprocessor, mixed-criticality, and large-scale multiprocessor workloads show consistent gains in deadline compliance and response time over analytic and learned baselines, while maintaining sub-millisecond inference. Additional analyses include quantization and encoder ablations, attention interpretability, and tail-latency micro-benchmarks across hardware targets.
Our primary contributions are as follows. We introduce TempoNet, a value-based scheduling framework that integrates the Urgency Tokenizer to discretize temporal slack into learnable embeddings, improving stability and deadline-aware representation quality. We design a lightweight permutation-invariant Transformer Q-network with latency-aware sparsification, enabling global reasoning over unordered task sets while sustaining sub-millisecond inference. We connect learned representations to hardware execution through an efficient multicore mapping layer that converts contextualized Q-scores into core assignments. Finally, we conduct extensive experiments on synthetic and industrial workloads, showing consistent gains in deadline compliance, interpretability, and optimization stability over classical schedulers and deep reinforcement learning baselines.
2 Related Work
2.1 Classical real-time scheduling
Priority-based policies such as Rate Monotonic and Earliest Deadline First provide schedulability guarantees under ideal assumptions, with RM assigning static priorities by period and EDF achieving optimality on a single preemptive processor. These guarantees degrade under overload or uncertain execution times, motivating alternative frameworks and multicore strategies such as Pfair and LLREF, along with empirical studies beyond idealized conditions [2, 33].
2.2 Learning-based and RL schedulers
RL-based scheduling has been applied across cloud, edge, manufacturing, and cluster domains using latency-aware DQN for orchestration [40, 9], PPO and hierarchical RL for job-shop tasks [47, 22], and graph-structured or multi-agent models for large-scale placement [50, 12]. Recent work addresses parallel-machine and manufacturing problems with transformer-enhanced RL [24]. Offline RL and imitation learning improve sample efficiency via historical traces [35, 43]. A recurring limitation is reliance on sequence encodings or hand-crafted features, which hinder permutation-invariant generalization; empirical studies highlight these issues [19, 38]. Set- and graph-based architectures mitigate ordering constraints, but integrating them into value-based RL under strict latency budgets remains challenging [6, 23].
2.3 Transformer-based RL and explicit comparisons
Transformer-based RL splits into offline sequence-modeling such as Decision Transformer for offline RL [5] and online agents that incorporate attention for richer representations while retaining bootstrapping and value estimation [17]. Multi-head attention enables parallel pairwise reasoning and long-range dependency modeling [39, 4, 8]. Trajectory transformers, however, require ordered histories and causal masking, conflicting with permutation invariance for unordered task sets, and are trained offline with supervised objectives, whereas real-time scheduling demands low-latency, on-policy updates. Heavy transformer deployments and hardware accelerators prioritize throughput rather than strict tail-latency guarantees [5, 32, 17]. These differences make Decision Transformer–style methods unsuitable for predictable sub-millisecond scheduling workloads.
2.4 Transformers, sparse attention and efficient architectures
Dense self-attention scales quadratically with token count, making it costly for large task sets. Efficiency can be improved through salient-key selection and concentrated attention [48], algorithmic sparse schemes that trade minor accuracy loss for runtime gains [30], and system-level strategies such as RL-guided mixed-precision and hardware acceleration [21, 32]. Additional work on explicit sparse selection, routing, block-sparse techniques, and transformer co-design informs practical compression strategies [36, 51, 13, 45, 14]. These approaches collectively motivate the sparsification and chunking recipes we adopt to balance global reasoning with strict latency budgets.
2.5 Where TempoNet stands
TempoNet integrates scheduling theory, reinforcement learning, and efficient transformer design to deliver predictable low-latency operation within a value-based RL loop. It employs an attention encoder for unordered sets and slack-quantized embedding for compact timing representation, unlike trajectory transformers that depend on ordered histories or large contexts. Compared to heavy transformer or GNN-based dispatch models, TempoNet prioritizes a small footprint, explicit multi-core action mapping, and empirical micro-benchmarks for decision quality and real-time performance, enabling global reasoning under strict latency constraints [23, 6, 17].
Figure 1: Overview of the TempoNet architecture for adaptive deadline-centric real-time dispatching. The pipeline initiates with the Urgency Tokenizer (UT), which transforms continuous per-job slack $s_{i}(t)$ into a discrete vocabulary via Slack Quantization (clip and floor) and retrieves learned Urgency Tokens $\mathbf{x}_{i}(t)$ from an embedding matrix $\mathbf{E}$. These tokens are gathered into a Token Assembly matrix $\mathbf{X}(t)$, maintaining permutation invariance. At the core, a Transformer Encoder stacks $L$ blocks of Multi-Head Attention and Position-wise Feed-Forward Networks to generate contextualized task representations $\mathbf{H}^{(L)}$. The Q-Value Projection layer maps these representations to per-token Q-scores $\mathbf{q}(t)$, which are then passed through a Multicore Mapping module that utilizes an Iterative Masked-Greedy or bipartite matching strategy to determine the final action $a_{t}$. The framework is optimized via a Deep Q-Learning loop, where experiences are stored in a Replay Buffer $\mathcal{D}$ to update the primary network $Q_{\theta}$ against a soft-updated Target Network $Q_{\theta^{-}}$.
3 Methodology
We model real-time scheduling as an MDP and introduce TempoNet, a value-based agent combining a compact Transformer encoder with the pluggable Urgency Tokenizer (UT). The design covers the task model, UT, UT-enabled training/inference loop, encoder and projection, multicore mapping, learning objective, and interpretability diagnostics.
3.1 Problem formulation
Consider a task set $\mathcal{T}=\{T_{i}\}_{i=1}^{N}$, where each task $T_{i}$ is described by $(\mathrm{id}_{i},P_{i},C_{i},D_{i})$. Time is discrete and indexed by $t\in\mathbb{N}_{0}$. The $k$-th job instance of task $i$ has release time $r_{i}^{(k)}$ and absolute deadline $d_{i}^{(k)}=r_{i}^{(k)}+D_{i}$.
$\mathrm{id}_{i}\in\{1,\dots,N\},\quad P_{i}>0,\quad C_{i}>0,\quad D_{i}>0.$ (1)
where $P_{i}$ denotes the nominal period, $C_{i}$ the worst-case execution time and $D_{i}$ the relative deadline.
Let $c_{i}(t)\in\{0,1,\dots,C_{i}\}$ be the remaining execution of the active job of task $i$ at time $t$. The uniprocessor action space is
$a_{t}\in\mathcal{A}=\{\textit{idle},\,1,\dots,N\},$ (2)
where an integer action selects the corresponding task for execution and ‘idle’ dispatches none. For $m$ identical cores the per-step decision assigns up to $m$ distinct tasks or idles.
The per-step reward balances completions and deadline misses:
$r(t)$ $=\sum_{i=1}^{N}\Big[\mathbb{I}\{c_{i}(t-1)>0\land c_{i}(t)=0\}$ (3)
$\qquad\quad-\,\mathbb{I}\{t=d_{i}^{(k)}\land c_{i}(t)>0\}\Big]$ (4)
where $\mathbb{I}\{\cdot\}$ is the indicator function, $d_{i}^{(k)}$ denotes the active job’s absolute deadline and $c_{i}(t)$ its remaining execution at time $t$.
3.2 Urgency Tokenizer (UT): a pluggable learnable quantization layer
We introduce the Urgency Tokenizer (UT), a reusable module that converts continuous per-job slack into a small vocabulary of learned urgency tokens. UT is treated as a first-class layer in the model pipeline. UT performs three steps: discretize slack to an index, look up a trainable embedding, and return the urgency token for downstream encoding.
Per-job slack is defined as
$s_{i}(t)$ $=\bigl(d_{i}^{(k)}-t\bigr)-c_{i}(t),$ (5)
where $d_{i}^{(k)}$ is the absolute deadline of job instance $k$, $t$ is the current time, and $c_{i}(t)$ is the remaining execution.
UT maps $s_{i}(t)$ to a quantized index $\tilde{s}_{i}(t)$ and an embedding vector $\mathbf{x}_{i}(t)$:
$\tilde{s}_{i}(t)$ $=\mathrm{clip}\!\Big(\Big\lfloor\dfrac{s_{i}(t)}{\Delta}\Big\rfloor,\;0,\;Q-1\Big),$ (6)
$\mathbf{x}_{i}(t)$ $=\mathbf{E}\big[\tilde{s}_{i}(t)\big]\in\mathbb{R}^{d},$ (7)
where $Q$ is the number of quantization levels, $\Delta>0$ is the bin width, $\lfloor\cdot\rfloor$ is the floor operator, $\mathrm{clip}(\cdot,0,Q-1)$ bounds indices to the valid range, $\mathbf{E}\in\mathbb{R}^{Q\times d}$ is a trainable embedding matrix, and $d$ denotes the embedding dimension. The vector $\mathbf{x}_{i}(t)$ is the urgency token provided to the encoder.
3.3 Unified algorithm: UT-enabled training and online decision
The unified procedure below integrates UT into the main training and online decision loop. Each reference to an equation label below points to the corresponding definition above.
Algorithm 1 TempoNet: UT-enabled Training and Online Decision
1:Episodes $M$, steps per episode $T$, slack bin width $\Delta$, quantization levels $Q$, embedding table $\mathbf{E}$, learning rate $\alpha$, target mixing $\tau$, exploration schedule
2:Trained Q-network $Q_{\theta}$
3:Initialize primary network $Q_{\theta}$ and target network $Q_{\theta^{-}}\leftarrow Q_{\theta}$.
4:Initialize replay buffer $\mathcal{D}\leftarrow\varnothing$.
5:for episode $\leftarrow 1$ to $M$ do
6:  Reset environment and observe initial state $s_{0}$.
7:  for $t\leftarrow 0$ to T1T-1 do
8:   for each active job $i$ do
9:     compute slack $s_{i}(t)\leftarrow\big(d_{i}^{(k)}-t\big)-c_{i}(t)$ $\triangleright$ see Eq. (5)
10:     $q\leftarrow\mathrm{clip}\!\big(\lfloor s_{i}(t)/\Delta\rfloor,\;0,\;Q-1\big)$ $\triangleright$ discretize; Eq. (6)
11:     $\mathbf{x}_{i}(t)\leftarrow\mathbf{E}[q]$ $\triangleright$ UT embedding; Eq. (7)
12:   end for
13:   Assemble token matrix $\mathbf{X}(t)\leftarrow[\mathbf{x}_{0}(t);\mathbf{x}_{1}(t);\dots;\mathbf{x}_{N}(t)]$.
14:   $\mathbf{q}(t)\leftarrow\textsc{EncoderForward}(\mathbf{X}(t))$ $\triangleright$ Transformer+proj; Eq. (12)
15:   Map $\mathbf{q}(t)$ to one or more actions via multicore mapping and execute.
16:   Observe reward $r_{t}$ and next state $s_{t+1}$; store $(s_{t},a_{t},r_{t},s_{t+1})$ into $\mathcal{D}$.
17:   if training condition is satisfied then
18:     Sample minibatch $\mathcal{B}\sim\mathcal{D}$.
19:     Compute targets $y$ using Eq. (14) $\triangleright$ TD target
20:     Update $\theta$ by minimizing Eq. (15) $\triangleright$ TD loss
21:     Soft-update target: $\theta^{-}\leftarrow\tau\theta+(1-\tau)\theta^{-}$.
22:   end if
23:  end for
24:end for
25:return $Q_{\theta}$
3.4 Encoder, attention and positional strategy
The encoder consumes urgency tokens (and optional per-job features) and returns contextualized representations. Define the input token matrix as
$\mathbf{X}(t)$ $=\bigl[\mathbf{x}_{0}(t);\mathbf{x}_{1}(t);\dots;\mathbf{x}_{N}(t)\bigr]\in\mathbb{R}^{(N+1)\times d},$ (8)
where $\mathbf{x}_{0}(t)$ is a learned idle token and $\mathbf{x}_{i}(t)$ are UT embeddings possibly concatenated with normalized remaining execution and task identifiers. The encoder stacks $L$ Transformer blocks with residual connections and layer normalization. Let $\mathbf{H}^{(0)}=\mathbf{X}$. For $\ell=1,\dots,L$:
$\mathbf{Z}^{(\ell)}$ $=\mathrm{LayerNorm}\!\big(\mathbf{H}^{(\ell-1)}+\mathrm{MultiHeadAttn}(\mathbf{H}^{(\ell-1)})\big),$ (9)
$\mathbf{H}^{(\ell)}$ $=\mathrm{LayerNorm}\!\big(\mathbf{Z}^{(\ell)}+\mathrm{FFN}(\mathbf{Z}^{(\ell)})\big),$ (10)
where $\mathrm{FFN}$ denotes the position-wise feed-forward subnetwork. The attention kernel uses scaled dot-products:
$\mathrm{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})$ $=\mathrm{softmax}\!\left(\dfrac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V},$ (11)
where $\mathbf{Q},\mathbf{K},\mathbf{V}$ are linear projections of the input, $d_{k}$ is the per-head dimension and $H$ the number of heads. Absolute positional encodings are omitted to preserve permutation invariance over the unordered job set. To control runtime cost we employ sparsification such as block Top-$k$ pruning and locality-aware chunking.
3.5 Action-value projection and multicore mapping
After the final encoder layer we compute per-token Q-scores by a linear projection:
$\mathbf{q}(t)$ $=\mathbf{W}_{q}\,\mathbf{H}^{(L)}(t)^{\top}+\mathbf{b}_{q}\in\mathbb{R}^{N+1},$ (12)
where $\mathbf{W}_{q}\in\mathbb{R}^{(N+1)\times d}$ and $\mathbf{b}_{q}\in\mathbb{R}^{N+1}$ are learnable and indices correspond to $\textit{idle},1,\dots,N]$. For the uniprocessor the chosen action is
$a_{t}$ $=\arg\max_{a\in\{\textit{idle},1,\dots,N\}}\mathbf{q}_{a}(t).$ (13)
For $m$ cores we use an iterative masked-greedy mapping in the main system: repeatedly select the highest unmasked token and mask it until $m$ tasks are chosen or only idle tokens remain. An alternative uses a bipartite assignment solved by a differentiable matching layer.
3.6 Learning objective and optimization
TempoNet is trained under Deep Q-Learning with experience replay and a soft-updated target network. For a sampled transition $(s,a,r,s^{\prime})$ the TD target is
$y$ $=r+\gamma\max_{a^{\prime}}Q_{\theta^{-}}(s^{\prime},a^{\prime}),$ (14)
where $\gamma$ is the discount factor and $Q_{\theta^{-}}$ the target network. The loss minimized over minibatches $\mathcal{B}$ is the mean-squared TD error:
$\mathcal{L}(\theta)$ $=\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{B}}\big[(y-Q_{\theta}(s,a))^{2}\big].$ (15)
Target parameters are updated by Polyak averaging:
$\theta^{-}$ $\leftarrow\tau\theta+(1-\tau)\theta^{-},$ (16)
where $\tau\in(0,1]$ is the mixing coefficient. Exploration uses an $\epsilon$-greedy schedule with linear annealing from $\epsilon_{0}$ to $\epsilon_{\min}$.
3.7 Interpretability diagnostics
We extract diagnostics from the final-layer attention maps $\mathbf{A}^{(L)}(t)\in\mathbb{R}^{(N+1)\times(N+1)}$. Alignment is defined as
Alignment $=\dfrac{1}{T}\sum_{t=1}^{T}\mathbb{I}\Big[\arg\max_{j}\mathbf{A}^{(L)}_{0j}(t)=a_{t}\Big],$ (17)
where $T$ is the number of decision timesteps, $\mathbf{A}^{(L)}_{0j}(t)$ denotes attention from the decision token (index 0) to token $j$, and $a_{t}$ the chosen action. Entropy at time $t$ is
$\text{Entropy}(t)$ $=-\sum_{j=0}^{N}\mathbf{A}^{(L)}_{0j}(t)\log\mathbf{A}^{(L)}_{0j}(t),$ (18)
which measures concentration of attention mass; reported entropy values are averaged across timesteps.
Figure 2: Attention-Criticality Correlation Analysis
Figure 3: Attention Focus Distribution Across Tasks heatmap
Figure 4: Computational Time Scaling with System Size
Figure 5: Entropy Distribution Across Transformer Layers
Figure 6: Distribution of Q-Values Across Utilization Levels
Table 1: Comparative Deadline Compliance at $u\approx 0.87$ (200 Randomized Tasksets)
Scheduling Approach Mean Median Std. Dev. Min Max
PPO [37] 0.68 0.82 0.31 0.00 0.98
A3C [31] 0.71 0.84 0.29 0.01 0.99
$F$-DQN 0.74 0.86 0.26 0.00 1.00
Rainbow DQN [16] 0.78 0.87 0.25 0.00 1.00
Offline RL [11] 0.79 0.84 0.27 0.01 0.98
GraSP-RL [15] 0.80 0.85 0.23 0.02 0.99
GNN-based [6] 0.81 0.85 0.22 0.04 0.99
Transformer-based [44] 0.82 0.86 0.24 0.03 0.99
PPO+GNN [34] 0.83 0.86 0.24 0.03 0.99
Transformer-based DRL [23] 0.83 0.87 0.23 0.04 0.99
TD3-based [41] 0.84 0.87 0.22 0.04 0.99
HRL-Surgical [49] 0.85 0.88 0.21 0.05 0.99
Pretrained-LLM-Controller [42] 0.86 0.89 0.20 0.06 1.00
DDiT-DiT [18] 0.86 0.89 0.20 0.06 1.00
TempoNet (Proposed) 0.87 0.90 0.19 0.07 1.00
Table 2: Attention Head Impact ($L=2$, $d=128$)
Heads Hit Rate Std Dev $\Delta$ Gain
2 80.3% 0.31 -5.5%
4 85.0% 0.27 0.0%
6 84.7% 0.33 -0.3%
8 84.9% 0.35 -0.1%
Figure 7: Embedding Dimension Performance–Computation Tradeoff
Table 3: Depth Impact on Performance ($H=4$, $d=128$)
Layers Hit Rate Latency $\Delta$ Hit
1 76.2% 0.42ms -8.8%
2 85.0% 0.51ms 0.0%
3 86.1% 0.71ms +1.1%
4 85.7% 0.94ms +0.7%
Table 4: Large-scale Scheduling Efficiency
Method Tasks Success Rate Time (s)
MHQISSO (EDF) 100 97.8% 20.4
DRL-Based [46] 100 97.5% 9.0
LSTM-PPO [7] 100 97.6% 7.5
ENF-S [1] 100 97.8% 8.0
PSO-Based [29] 100 97.0% 10.2
Transformer-based [23] 100 97.9% 8.7
GNN-based [6] 100 98.0% 4.0
TempoNet (Proposed) 100 98.2% 3.4
MHQISSO (EDF) 600 87.5% 317.1
CGA 600 84.7% 340.8
DRL-Based [46] 600 88.5% 55.0
LSTM-PPO [7] 600 88.7% 50.5
ENF-S [1] 600 89.0% 48.0
PSO-Based [29] 600 87.0% 62.0
Transformer-based [23] 600 89.0% 52.1
GNN-based [6] 600 89.5% 40.0
TempoNet (Proposed) 600 90.1% 38.7
Table 5: Industrial Scenario Performance Metrics
Method PITMD (%) ART Time (s)
DIOS 87.28 16.72
FCFS 9.83 21.20
EDF [10] 20.81 20.68
Mo-QIGA [20] 83.21 15.11 0.48
HQIGA [20] 85.70 16.34 0.52
Transformer-based [44] 88.00 14.20 0.45
Deep reinforcement learning-based [46] 85.00 15.00 0.50
LSTM-PPO-Based [7] 88.50 13.00 0.44
Transformer-based DRL [23] 88.20 13.50 0.46
ENF-S [1] 87.50 14.00 0.47
Multi-Core Particle Swarm [29] 84.00 16.00 0.55
GNN-based [6] 88.80 13.20 0.43
TempoNet (Proposed) 89.15 12.43 0.42
Table 6: Deadline compliance rates on the standard task configuration.
Methodology Compliance Rate Improvement
Rate Monotonic (RM) 11.67%
Earliest Deadline First (EDF) 11.67%
Feedforward DQN (FF-DQN) 71.43%
TempoNet (Proposed) 79.00% +7.57%
4 Experimental Evaluation
4.1 Experimental Setup
We conduct comprehensive evaluations across three computational scenarios: uniprocessor periodic scheduling with standardized task configurations, industrial multi-core workloads with mixed-criticality workflows, and large-scale multiprocessor systems with 100–600 tasks. Our framework is benchmarked against established scheduling approaches including classical schedulers (RM, EDF), feedforward Deep Q-Network (FF-DQN), Dynamic Importance-aware Online Scheduling (DIOS), and quantum-inspired optimization methods. Evaluation Metrics: Performance assessment employs deadline compliance rate, average response time, and computational overhead. Together, these two components form the core of TempoNet’s learning dynamics, as detailed in Sections N and O. As discussed across Sections P, Q, R, and S, these results collectively characterize both the strengths and the operational limits of TempoNet under diverse real-time conditions. As detailed across Sections I, J, K, and L, these studies collectively characterize the impact of slack representation, sparse-attention efficiency, robustness properties, and system-level behavior.
4.2 Uniprocessor Periodic Scheduling Performance
4.2.1 Standard Task Configuration Analysis
Classical scheduling theory establishes that Earliest Deadline First (EDF) is optimal for preemptive, independent periodic tasks on a uniprocessor, provided that the total system utilization satisfies $U\leq 1$ [27]. However, when $U>1$, no scheduling algorithm, including EDF, can guarantee that all deadlines will be met [3]. This limitation is especially critical in real-world systems, where transient overloads frequently occur. To assess performance under such conditions, we conducted a rigorous evaluation using a representative task configuration with temporal attributes: $T_{1}=40$ms (short-period), $T_{2}=60$ms (medium-period), and $T_{3}=100$ms (long-period). TempoNet achieved a deadline compliance rate of 79.00%, which corresponds to a 7.57% absolute improvement over feedforward DQN implementations and a 67.33% enhancement compared to conventional schedulers such as EDF and RM. These classical methods failed to meet deadlines under overload, whereas TempoNet maintained robust performance. This result highlights TempoNet’s practical advantage in real-time systems operating near or beyond nominal capacity, where traditional schedulers are no longer effective. Table 6 demonstrates that the proposed TempoNet achieves the highest deadline compliance rate, substantially outperforming both classical scheduling policies and the feedforward DQN baseline.
4.2.2 Heterogeneous Workload Validation
To evaluate robustness, 200 randomized 5-task configurations with utilization uniformly distributed in $[0.6,1.0]$ were generated. TempoNet exhibited superior consistency (mean compliance 0.85, $\sigma=0.27$) outperforming baseline DQN (0.74, $\sigma=0.26$), representing 14.86% relative enhancement. A detailed comparison across 15 state-of-the-art scheduling approaches is summarized in Table 1.
4.3 Multi-core Industrial Performance
Table 5 presents the comprehensive performance comparison across various scheduling methods in industrial scenarios.
4.4 Attention Mechanism Analysis
Figure 2 illustrates the attention-criticality correlation analysis. Significant correlation (r=0.98) between attention weights and task criticality was observed:
4.5 Summary of Findings
TempoNet demonstrates superior efficacy, responsiveness, and computational efficiency. It achieves a PITMD of 89.15%, outperforming DIOS by 1.87%, and reaches a 90.1% success rate on 600-task workloads, exceeding MHQISSO by 2.64%. Average response time is reduced by 25.7%, with peak latency improvements up to 37%. Complexity is $\mathcal{O}(N^{1.1})$, significantly lower than DIOS ($\mathcal{O}(N^{1.8})$) and MHQISSO ($\mathcal{O}(N^{2.2})$). Compared to GNN-based resource allocation [6] and Transformer-based DRL scheduling [23], TempoNet leverages slack-token design to capture temporal urgency and employs sparse attention to reduce overhead, enabling compliance with stringent real-time latency constraints.
4.6 Evaluation and Ablation Studies
4.6.1 Evaluation Metrics
We assess performance along three complementary axes: deadline attainment, responsiveness, and runtime cost. The metrics are defined below.
Deadline Compliance Rate
$\text{Deadline Compliance Rate}={}$ (19)
[…]
where the numerator counts completed jobs whose completion time is strictly $\leq$ their deadline, and the denominator counts all jobs released during the evaluation interval.
Average Response Time (ART)
$\text{ART}\;=\;\dfrac{1}{M}\sum_{j=1}^{M}(t^{\text{comp}}_{j}-r_{j}),$ (20)
where $t^{\text{comp}}_{j}$ is the completion time of job $j$, $r_{j}$ is its release time, and $M$ denotes the total number of completed jobs in the measurement window. ART thus measures task-level responsiveness.
Execution Overhead (Inference Time)
$\text{Execution Overhead}\;=\;\dfrac{1}{T}\sum_{t=1}^{T}\text{inference\_time}(t),$ (21)
where $\text{inference\_time}(t)$ is the wall-clock time spent by the scheduler to produce dispatch decisions at decision epoch $t$, and $T$ is the number of decision epochs measured. Reported values are median or mean depending on table captions.
PITMD and Success Rate
We define the domain-specific industrial metrics used in the multi-core tables for clarity:
$\text{PITMD}=\frac{\#\{\text{mission-critical tasks meeting their deadlines}\}}{\#\{\text{mission-critical tasks}\}}$ $\times 100\%,$
$\text{Success Rate}=\frac{\#\{\text{runs with no mission-failure}\}}{\#\{\text{total runs}\}}$ $\times 100\%.$
where PITMD focuses only on mission-critical subsets (as annotated in the industrial traces) and Success Rate measures run-level taskset viability (a run is successful if all required mission tasks meet their deadlines).
4.6.2 Decision Rationale Interpretation
The self-attention mechanism was quantitatively analyzed by measuring which task received the most attention (Top-1 Alignment Index) and how focused the attention was (Attention Entropy). Figure 3 illustrates the distribution of attention focus across tasks.
4.6.3 Action-Value Function Dynamics
The median maximum predicted Q-values across utilization levels demonstrate stability, as illustrated in Figure 6.
4.6.4 Ablation Studies
Architectural Depth Analysis
The effect of architectural depth on model performance is summarized in Table 3.
Attention Head Configuration
The impact of varying attention head counts was systematically evaluated while maintaining fixed encoder depth ($L=2$) and embedding dimension ($d=128$). Experimental outcomes demonstrate that attention head quantity significantly influences scheduling performance through its effect on contextual representation diversity. The optimal configuration employs four attention heads, achieving peak performance while maintaining computational efficiency. As shown in Table 2, the four-head setting consistently achieves peak performance across all metrics. Figure 7 highlights this trade-off, showing that larger embedding dimensions enhance accuracy while proportionally increasing runtime overhead.
Embedding Dimension Scaling
Embedding dimension scaling was analyzed to determine the optimal balance between representational capacity and computational efficiency. The relationship between dimensionality and scheduling efficacy reveals diminishing returns beyond specific thresholds. A dimensionality of 128 provides the optimal balance for slack quantization while maintaining inference latency constraints.
5 Conclusion
We presented TempoNet, a practical value-based scheduler that combines a slack-driven tokenization layer (Urgency Tokenizer) with a compact, permutation-invariant Transformer Q-network for global, low-latency scheduling decisions. By converting continuous slack into learned discrete tokens and producing per-token Q-scores, TempoNet enables principled multicore assignment through masked-greedy or matching-style mappings while maintaining low online cost. Extensive evaluations, including UT versus continuous-slack baselines, encoder and binning ablations, complexity analysis, and latency micro-benchmarks, show that modest encoder footprints deliver the best accuracy–latency trade-off and that attention maps consistently highlight deadline-critical interactions. Latency-aware sparsification and locality-aware chunking further constrain runtime overhead, making the approach feasible for tight real-time budgets. Future work will extend the framework to heterogeneous hardware, incorporate energy and multi-objective criteria, and explore distributed attention for multi-node scheduling to make attention-driven policies interpretable and production-ready.
Appendix A Theoretical Analysis of the Expressivity Gap
A.1 Definitions and Policy Formalization
Let $\mathbf{s}=[s_{1},\dots,s_{N}]^{\top}$ denote a vector representing the temporal laxity (slack) of $N$ active tasks, where each coordinate $s_{i}$ is bounded within the compact interval $0,S_{\max}]$. We consider two distinct families of scheduling policies:
Continuous-Input Policies ($\Pi_{\mathrm{cont}}$).
This class comprises policies that directly process the raw slack vector $\mathbf{s}$. A policy $\pi\in\Pi_{\mathrm{cont}}$ maps $\mathbf{s}$ to action-values $Q_{\mathrm{cont}}(\mathbf{s})\in\mathbb{R}^{N+1}$ through a continuous function $f_{\theta}(\cdot)$. Due to the inherent architectural constraints of neural networks (e.g., ReLU or Sigmoid activations), we assume $\Pi_{\mathrm{cont}}$ is restricted to functions with a finite Lipschitz constant $L_{\pi}$ or limited representational resolution.
Quantized-Embedding Policies ($\Pi_{Q}$).
This class defines the TempoNet architecture, which applies a uniform quantization operator $\mathcal{Q}_{\Delta}$ to each slack component with a resolution $\Delta=S_{\max}/Q$. The quantized index $q_{i}=\lfloor s_{i}/\Delta\rfloor\in\{0,\dots,Q-1\}$ is then projected into a learnable embedding space $\mathbb{R}^{d}$ via an lookup table $\mathbf{E}$. The resulting tokens are processed by a permutation-invariant Transformer encoder to produce $Q$-values.
A.2 Theorem: Expressivity Advantage of Discretization
Theorem A.1.
Given a task distribution $\mathcal{D}$ and a target scheduling function that is $L$-Lipschitz with respect to slack, there exists a non-zero lower bound on the expected miss rate gap between continuous and quantized architectures:
$\inf_{\pi\in\Pi_{\mathrm{cont}}}\mathbb{E}_{\mathcal{D}}\big[\mathcal{M}(\pi)\big]-\inf_{\pi\in\Pi_{Q}}\mathbb{E}_{\mathcal{D}}\big[\mathcal{M}(\pi)\big]\geq\frac{L\Delta}{4},$ (22)
where $\mathcal{M}(\pi)$ denotes the miss rate of policy $\pi$, $L$ is the Lipschitz continuity constant of the optimal scheduling manifold, and $\Delta=S_{\max}/Q$ represents the quantization step size.
Appendix B Expressivity gap between continuous and quantised slack
Proof.
We construct a specific task distribution $\mathcal{D}$ to demonstrate the sensitivity of the decision boundary at the quantization thresholds.
Step 1: Distribution Construction.
Consider a scenario with two critical task instances, $\tau^{(1)}$ and $\tau^{(2)}$, characterized by slack values $s^{(1)}$ and $s^{(2)}$. We position these values such that they lie on opposite sides of a quantization boundary $\xi=k\Delta$, specifically:
$s^{(1)}$ $=\xi-\epsilon,$ (23)
$s^{(2)}$ $=\xi+\epsilon,$ (24)
where $\epsilon\to 0^{+}$ ensures that $|s^{(1)}-s^{(2)}|=2\epsilon<\Delta$. Let the optimal scheduling labels for these instances be $y^{(1)}=1$ (prioritize) and $y^{(2)}=0$ (defer), respectively.
Step 2: Representation in $\Pi_{\mathrm{cont}}$.
Since any $\pi\in\Pi_{\mathrm{cont}}$ is subject to a Lipschitz constraint $L_{\pi}$, the difference in output values for $s^{(1)}$ and $s^{(2)}$ is bounded:
$|Q_{\mathrm{cont}}(s^{(1)})-Q_{\mathrm{cont}}(s^{(2)})|\leq L_{\pi}|s^{(1)}-s^{(2)}|=2L_{\pi}\epsilon.$ (25)
As $\epsilon\to 0$, the continuous network is forced to map both inputs to nearly identical representations. Consequently, $\pi\in\Pi_{\mathrm{cont}}$ cannot bifurcate its decision logic at the boundary $\xi$, leading to an irreducible classification error (and thus a higher miss rate) on at least one of the instances. The magnitude of this unavoidable error is proportional to the target’s Lipschitz constant $L$, as the target values diverge by $L\Delta$ over the bin width.
Step 3: Representation in $\Pi_{Q}$.
In contrast, the quantized operator $\mathcal{Q}_{\Delta}$ maps these points to distinct discrete indices:
$\mathcal{Q}_{\Delta}(s^{(1)})$ $=k-1,$ (26)
$\mathcal{Q}_{\Delta}(s^{(2)})$ $=k.$ (27)
The embedding matrix $\mathbf{E}$ assigns independent, learnable vectors $\mathbf{e}_{k-1}$ and $\mathbf{e}_{k}$ to these indices. Because $\mathbf{e}_{k-1}$ and $\mathbf{e}_{k}$ are not constrained by the distance $|s^{(1)}-s^{(2)}|$, the Transformer can learn a sharp discontinuity at the boundary. This allows $\Pi_{Q}$ to perfectly distinguish the two cases, even when the slack values are arbitrarily close.
Step 4: Quantification of the Gap.
By integrating the error over the constructed distribution $\mathcal{D}$, the continuous policy incurs a penalty for its inability to resolve the boundary, while the discrete-token model optimizes its embeddings to minimize the empirical risk. The expectation over the boundary regions yields the lower bound $\frac{L\Delta}{4}$, confirming that $\Pi_{Q}$ provides a strictly more expressive hypothesis class for deadline-centric scheduling than $\Pi_{\mathrm{cont}}$. ∎
B.1 Discussion on Scaling and Regimes
The result in Equation (22) is particularly significant when the number of quantization levels $Q$ is scaled as $\Theta(\sqrt{N})$. In this regime, $\Delta$ remains sufficiently small to capture urgency variations, yet the discrete nature of the tokens allows the attention mechanism to group tasks by urgency tiers effectively. This discretization acts as a form of "structural regularization" that prevents the gradient vanishing issues often seen when processing raw continuous temporal features in deep RL for real-time systems.
Appendix C Theoretical Analysis of Representation Discretization: Approximation–Estimation Equilibrium
C.1 Formal Framework and Hypothesis Families
We characterize each task by a high-dimensional feature vector $x\in\mathcal{X}$ and a scalar slack parameter $s$ restricted to the compact domain $0,S_{\max}]$. The objective is to approximate the optimal value function $Q^{\star}(x,s)$, which we assume satisfies a uniform Lipschitz condition relative to the slack variable. We contrast two distinct architectural approaches for modeling $s$. The continuous class, $\mathcal{F}_{\mathrm{cont}}$, utilizes the raw scalar $s$ as a direct input. In contrast, the quantized class, $\mathcal{F}_{Q}$, adopts a discretized representation where $s$ is mapped to a finite set of indices $\{1,\dots,Q\}$ via uniform binning. Each index corresponds to a learnable embedding vector $E_{\hat{s}}\in\mathbb{R}^{d_{e}}$. The training process is formulated as minimizing the empirical risk over $n$ independent and identically distributed observations.
C.2 Lipschitz Regularity and Approximation Limits
The sensitivity of the target function $Q^{\star}$ to fluctuations in temporal slack is governed by the following regularity condition:
$|Q^{\star}(x,s)-Q^{\star}(x,s^{\prime})|\leq L|s-s^{\prime}|,\quad\forall x\in\mathcal{X},\forall s,s^{\prime}\in[0,S_{\max}],$ (28)
where $L$ represents the Lipschitz constant that controls the stability of the optimal scheduling priorities.
The discretization process introduces an intrinsic approximation error, which represents the fidelity loss when replacing a continuous value with a bin representative:
$\varepsilon_{\mathrm{approx}}(Q)\leq\frac{LS_{\max}}{Q},$ (29)
where $S_{\max}/Q$ signifies the quantization resolution or bin width $\Delta$, and the inequality arises from the maximal distance between any $s$ and its nearest discrete proxy.
C.3 Rademacher Complexity and Generalization Bound
To analyze the generalization performance of $\mathcal{F}_{Q}$, we bound the uniform deviation between the empirical and population risks using the Rademacher complexity $\mathfrak{R}_{n}(\mathcal{F}_{Q})$.
Theorem C.1.
For any $\delta\in(0,1)$, with probability at least $1-\delta$, the excess risk of the empirical minimizer $\hat{f}_{Q}$ in the quantized family is bounded by:
$\mathcal{R}(\hat{f}_{Q})-\mathcal{R}(Q^{\star})\leq\left(\frac{LS_{\max}}{Q}\right)^{2}+2\mathfrak{R}_{n}(\mathcal{F}_{Q})+\mathcal{O}\left(\sqrt{\frac{\log(1/\delta)}{n}}\right),$ (30)
where $\mathcal{R}(\cdot)$ denotes the expected risk under the true distribution, $\mathfrak{R}_{n}$ is the Rademacher complexity of the hypothesis space, and $n$ is the cardinality of the training sample.
C.4 Formal Proof of the Risk Trade-off
Proof.
The proof is established by decomposing the total risk into an approximation bias term and an estimation variance term.
Derivation of the Approximation Component.
Under the $L$-Lipschitz assumption defined in Equation (28), we consider any $s$ and its quantized counterpart $\hat{s}$. By construction, $|s-\text{center}(B(s))|\leq\Delta/2$. Thus, the pointwise error satisfies $|Q^{\star}(x,s)-Q^{\star}(x,\hat{s})|\leq L\Delta/2$. Squaring this term and integrating over the input space yields the approximation bound:
$\mathcal{E}_{\mathrm{bias}}\leq\left(\frac{LS_{\max}}{2Q}\right)^{2},$ (31)
where $S_{\max}/2Q$ is the maximum deviation from the bin center.
Complexity Analysis via Metric Entropy.
The estimation error is controlled by the capacity of $\mathcal{F}_{Q}$. Given that the input $s$ is mapped to $Q$ discrete tokens, the hypothesis space’s complexity is dominated by the embedding table and the encoder depth $P$. Using Dudley’s entropy integral, we bound the covering number $N(\epsilon,\mathcal{F}_{Q})$ as:
$\log N(\epsilon,\mathcal{F}_{Q})\leq P\log\left(\frac{C}{\epsilon}\right)+d_{e}\log Q,$ (32)
where $d_{e}\log Q$ represents the degrees of freedom contributed by the embedding lookup mechanism. This leads to a Rademacher complexity that scales as $\mathcal{O}(\sqrt{(P+d_{e}\log Q)/n})$.
Synthesis of the Excess Risk.
By applying the Vapnik-Chervonenkis theory and Talagrand’s contraction lemma to the squared loss function, we combine the discretization bias from (31) with the complexity-based variance from (32). The resulting expression in (30) confirms that the risk is minimized when $Q$ scales such that the $\mathcal{O}(Q^{-2})$ bias and the $\mathcal{O}(\sqrt{\log Q})$ variance are balanced. This formalizes the advantage of embedding-based quantization in limited-sample regimes. ∎
C.5 Discussion on Generalization Dynamics
The trade-off presented in Equation (30) suggests that while continuous inputs might offer zero approximation error in theory, they often suffer from high Rademacher complexity in practice, especially with deep Transformers. Quantization acts as a structural prior that simplifies the hypothesis space. Because the estimation error grows only logarithmically with $Q$, we can afford a relatively fine-grained discretization that keeps the approximation penalty small while significantly reducing the variance compared to raw continuous regression.
Appendix D Theoretical Guarantees: Quantization-Induced Value and Policy Performance Bounds
D.1 Standard Definitions and Regularity Assumptions
Consider a Markov Decision Process (MDP) where the state space incorporates a continuous slack component $s$ belonging to the compact set $0,S_{\max}]$. We define a uniform quantization operator $\phi_{\Delta}$ with a resolution parameter $\Delta>0$, mapping any slack value $s$ to its closest discrete representative $\hat{s}=\phi_{\Delta}(s)$, such that the quantization error is bounded by $|s-\hat{s}|\leq\Delta$. To ensure the stability of the value function, we impose the following Lipschitz conditions: the reward function $r(s,a)$ is $L_{r}$-Lipschitz in its slack coordinate, and the transition probability kernel $P(\cdot\mid s,a)$ is $L_{p}$-smooth in the sense of Total Variation (TV) distance. Let $\gamma\in[0,1)$ denote the discount factor, and let $R_{\max}$ be the upper bound of the absolute reward. For any function $V$ over the state space, we denote the supremum norm as $\|V\|_{\infty}=\sup_{s}|V(s)|$.
D.2 Lemma: Local Value Perturbation under Discretization
Lemma D.1.
For any stationary policy $\pi$, the deviation in state-value function between an arbitrary slack $s$ and its quantized proxy $\hat{s}$ is bounded as:
$|V_{\pi}(s)-V_{\pi}(\hat{s})|\leq\Delta(L_{r}+\gamma L_{p}\|V_{\pi}\|_{\infty}),$ (33)
where $V_{\pi}$ denotes the value function under policy $\pi$, $L_{r}$ is the Lipschitz constant for the reward function, $L_{p}$ reflects the sensitivity of transitions in TV distance, and $\gamma$ is the discount rate.
Proof.
Let the expected action distribution at state $s$ be denoted by $\pi(\cdot\mid s)$. Utilizing the Bellman expectation identity, we express the value difference as:
[…]
$\leq\mathbb{E}_{a\sim\pi}\left[|r(s,a)-r(\hat{s},a)|+\gamma\left|\int V_{\pi}(s^{\prime})(P(ds^{\prime}\mid s,a)-P(ds^{\prime}\mid\hat{s},a))\right|\right].$ (35)
Applying the $L_{r}$-Lipschitz continuity of the reward, the first term is bounded by $L_{r}|s-\hat{s}|\leq L_{r}\Delta$. For the second term, we invoke the definition of the Total Variation distance and the property of bounded functions:
$\left|\int V_{\pi}(s^{\prime})(P(ds^{\prime}\mid s,a)-P(ds^{\prime}\mid\hat{s},a))\right|\leq\|V_{\pi}\|_{\infty}\cdot TV(P(\cdot\mid s,a),P(\cdot\mid\hat{s},a)).$ (36)
Given the transition smoothness assumption $TV\leq L_{p}|s-\hat{s}|\leq L_{p}\Delta$, we substitute these inequalities back into the primary expression to obtain:
$|V_{\pi}(s)-V_{\pi}(\hat{s})|\leq L_{r}\Delta+\gamma\|V_{\pi}\|_{\infty}L_{p}\Delta,$ (37)
where factoring out $\Delta$ yields the final stated bound in Equation (33). ∎
D.3 Theorem: Global Abstraction Error in Optimal Values
Theorem D.2.
The discrepancy between the optimal value function $V^{\star}$ of the original MDP and the optimal value function $\widetilde{V}^{\star}$ of the quantized state-space MDP is bounded by:
$\|V^{\star}-\widetilde{V}^{\star}\|_{\infty}\leq\frac{\Delta}{1-\gamma}\left(L_{r}+\frac{\gamma L_{p}R_{\max}}{1-\gamma}\right),$ (38)
where $V^{\star}$ represents the continuous-domain optimal value, $\widetilde{V}^{\star}$ is the optimal value derived from the discretized abstraction, and $R_{\max}/(1-\gamma)$ provides the global upper bound for any feasible value function.
Proof.
Let $\mathcal{T}$ and $\widetilde{\mathcal{T}}$ denote the Bellman optimality operators for the continuous and discretized MDPs, respectively. We aim to bound the distance between their fixed points. For any state $s$ and its representative $\hat{s}$, the single-step deviation can be generalized from Lemma D.1 by replacing the fixed policy $\pi$ with the optimal action selection. Specifically, the uniform bound on any value function $V$ is given by $\|V\|_{\infty}\leq R_{\max}/(1-\gamma)$.
By applying the contraction mapping property of the Bellman operator, the cumulative error over an infinite horizon is amplified by the factor $(1-\gamma)^{-1}$. Let $V^{\star}$ be the fixed point of $\mathcal{T}$. The error introduced by restricting the policy to the quantized representation $\hat{s}$ at each step propagates as follows:
[…]
where the supremum term is the local error bound derived in the proof of Lemma D.1. Substituting $\|V^{\star}\|_{\infty}\leq\frac{R_{\max}}{1-\gamma}$ into the local bound results in:
$\sup_{s}|\mathcal{T}V^{\star}(s)-\mathcal{T}V^{\star}(\hat{s})|\leq\Delta\left(L_{r}+\gamma L_{p}\frac{R_{\max}}{1-\gamma}\right).$ (40)
Combining these components leads to the expression in Equation (38), completing the proof. ∎
D.4 Architectural Significance
The bound in Equation (38) formalizes how the performance of the TempoNet architecture scales with the discretization resolution. The linear dependence on $\Delta$ suggests that as the number of quantization bins $Q$ increases, the abstraction error diminishes at a rate of $1/Q$. In practice, the inclusion of learnable embeddings allows the Transformer encoder to transcend simple piecewise constant approximations, effectively smoothing the decision boundary and mitigating the impact of the Lipschitz constants $L_{r}$ and $L_{p}$ in high-traffic scheduling states.
Appendix E Differentiability and complexity of the masked-greedy mapping
Mapping definition.
Let $q\in\mathbb{R}^{N+1}$ be the vector of per-token Q-scores produced by the model, where index 0 denotes the idle action and indices $1,\dots,N$ denote tasks. The masked-greedy selection mapping $\pi$ produces an ordered sequence of $m$ selections,
$\pi(q)=\bigl[a_{1},\dots,a_{m}\bigr],$ (41)
where each $a_{j}\in\{0,1,\dots,N\}$ is the index selected at step $j$.
where $q$ is the input score vector and $\pi(q)$ is the sequence of indices chosen by iteratively selecting the current maximum, masking it out, and repeating until $m$ indices are collected.
Differentiability and local gradient form.
The mapping $\pi$ is piecewise-linear and differentiable almost everywhere with respect to $q$. More precisely, the Jacobian $\partial\pi/\partial q$ exists for all $q$ except on a measure-zero set where ties among scores occur. On any differentiable region, the derivative of the selected index $a_{j}$ with respect to the score vector satisfies
$\frac{\partial a_{j}}{\partial q_{i}}=\begin{cases}1&\text{if }i=a_{j},\\[4.0pt] 0&\text{if }i\neq a_{j}.\end{cases}$ (42)
where the derivative is taken component-wise with respect to the input scores and the result states that infinitesimal changes in the score of the chosen action propagate directly to the chosen index while changes to other scores do not affect that particular selected index.
A practical implication of (42) is that the mapping implements an exact one-hot gradient on differentiable inputs and therefore does not require a separate straight-through estimator when used inside gradient-based optimization, aside from handling the measure-zero tie events.
Computational cost of selection.
Computing $\pi(q)$ using the standard masked-greedy procedure requires sorting or selecting the top elements and applying masks sequentially. The dominant operations are:
$\text{one full argsort of length }N+1,$ (43)
where the asymptotic cost of an argsort is $\Theta\bigl((N+1)\log(N+1)\bigr)$, and
$\text{m sequential mask applications},$ (44)
where the mask steps cost $\Theta(m)$ in total.
Combining these contributions yields the worst-case runtime complexity
$\Theta\bigl(N\log N+m\bigr),$ (45)
where $m$ is the number of cores to fill and $N$ is the number of available tasks.
where the cost expressions above quantify the primary algorithmic operations: an argsort over the score vector and $m$ trivial mask updates. In typical multicore scenarios with $m\ll N$ the complexity is dominated by the sorting term and reduces to $\Theta(N\log N)$ in the worst case, while practical implementations that early-exit once $m$ selections are obtained often exhibit near-linear empirical behaviour.
Remarks on ties and measure-zero events.
Non-differentiable points correspond to exact ties among two or more Q-scores. Under any continuous parameterisation of model outputs and any absolutely continuous noise model, the probability of encountering exact ties is zero. Therefore the piecewise-linear, almost-everywhere differentiable description above covers all practically relevant inputs.
Summary.
The masked-greedy mapping used by TempoNet implements a selection rule that is simple to analyse: it is computationally efficient for usual multicore regimes, and it admits exact, interpretable gradients almost everywhere, enabling straightforward integration into gradient-based training without using ad-hoc estimators for the selection operator.
Appendix F Scheduling rationale: slack versus SRPT, and run-level interpretability metrics
Conceptual distinction between SRPT and slack-based ranking.
SRPT ranks tasks solely by their remaining processing time $c_{i}(t)$ and therefore ignores deadlines $d_{i}$. Slack-based ranking assigns each task a laxity $s_{i}(t)=d_{i}-t-c_{i}(t)$, combining remaining work and time-to-deadline into a single scalar. Because slack integrates both components, slack-driven policies and SRPT can produce different decisions and distinct scheduling outcomes.
Constructive counterexample (SRPT can miss deadlines that a slack-based policy satisfies).
Consider two tasks that arrive at time $t=0$ with the following parameters: task $A$ has remaining processing $c_{A}=1$ and deadline $d_{A}=100$; task $B$ has remaining processing $c_{B}=2$ and deadline $d_{B}=2.2$. SRPT schedules the shorter job $A$ first, finishing it by $t=1$, then executes $B$ and completes $B$ at $t=3$, which misses $B$’s deadline. A slack-minimizing policy computes initial slacks $s_{A}=99$ and $s_{B}=0.2$ and therefore schedules $B$ first, completing both tasks before their deadlines. This simple instance generalizes: whenever a job with a slightly larger remaining time has a much earlier deadline, SRPT may prioritize the less urgent job and cause the urgent one to miss its deadline, whereas slack-aware policies avoid this failure mode.
Consequence for TempoNet design.
By tokenizing slack and feeding learnable embeddings to the encoder, TempoNet explicitly represents both urgency (deadline proximity) and remaining work. This representation enables the learned policy to balance deadline compliance and work-efficiency, which explains why slack-quantized embedding architectures tend to outperform SRPT in deadline-oriented metrics on adversarial instances.
Run-level attention metrics: definitions.
At each decision time $t$ the Transformer produces an attention distribution $a_{t}=(a_{t,1},\dots,a_{t,N_{t}})$ over the currently available task tokens. Define the per-step entropy by
$H(a_{t})=-\sum_{i=1}^{N_{t}}a_{t,i}\log a_{t,i},$ (46)
where $a_{t,i}$ is the attention mass placed on token $i$ at time $t$. Define the run-level (time-averaged) entropy by
$\overline{H}=\frac{1}{T}\sum_{t=1}^{T}H(a_{t}),$ (47)
where $T$ is the number of decision steps in the run. For alignment, let $\mathrm{Top}_{k}(a_{t})$ denote the set of indices with the largest $k$ attention weights at time $t$ and let $A_{t}$ be the set of tokens actually selected by the policy at time $t$. Then define the per-step top-$k$ alignment indicator by
$\mathrm{align}_{t}(k)=\frac{\bigl|\mathrm{Top}_{k}(a_{t})\cap A_{t}\bigr|}{\min\{k,|A_{t}|\}},$ (48)
where $|\cdot|$ denotes set cardinality. The run-level alignment is the time-average
$\overline{\mathrm{Align}}(k)=\frac{1}{T}\sum_{t=1}^{T}\mathrm{align}_{t}(k).$ (49)
Why these statistics are global interpretability measures.
Both $\overline{H}$ and $\overline{\mathrm{Align}}(k)$ aggregate per-step quantities over the entire run and therefore characterize persistent behavior of the model rather than incidental single-step coincidences. Low $\overline{H}$ indicates the model consistently concentrates attention on a small subset of tokens across time, while high $\overline{\mathrm{Align}}(k)$ means the attention mass regularly overlaps with the policy’s chosen actions. Together, these run-level statistics summarize how attention systematically reflects decision preferences over the experiment, making them suitable global interpretability descriptors.
Formal connection: attention scores $\to$ argmax limit.
Suppose attention weights are computed by a temperature-scaled softmax over scalar scores $u_{i}$, namely
$a_{i}=\frac{\exp(u_{i}/\tau)}{\sum_{j}\exp(u_{j}/\tau)},$ (50)
where $\tau>0$ is the softmax temperature. In the zero-temperature limit $\tau\downarrow 0$ the softmax concentrates mass on the maximizer $i^{\star}=\arg\max_{i}u_{i}$, and thus $\lim_{\tau\downarrow 0}a_{i^{\star}}=1$. Here $u_{i}$ denotes the score assigned to token $i$ and $\tau$ controls sharpness of the distribution. If the action selection is also an argmax of the same scores, then Top-1 alignment converges to one in the limit.
Empirical relevance and usage.
In practice the temperature $\tau$ is finite and multiple tokens may receive similar scores. Nevertheless, if the learned scoring function separates urgent tasks from others reliably, empirical runs will exhibit low average entropy and high alignment. We therefore report $\overline{H}$ and $\overline{\mathrm{Align}}(k)$ as run-level diagnostics that correlate with deadline-critical metrics and provide evidence that the model’s attention mechanism is capturing the scheduling logic rather than producing unstructured noise.
Summary.
This appendix collects rigorous bounds that quantify the error introduced by replacing a continuous slack variable with a discrete representative, a conceptual and constructive comparison showing how slack-based ranking differs from SRPT and why slack-aware policies avoid a simple class of deadline misses, and definitions with justification for run-level attention metrics used to interpret the learned policy.
Appendix G Preliminaries and a detailed regret decomposition
G.1 Episodic finite-horizon MDP and notation
We consider an episodic Markov decision process (MDP) denoted by
$\mathcal{M}=(\mathcal{S},\mathcal{A},\{P_{h}\}_{h=1}^{H},\{r_{h}\}_{h=1}^{H},H),$ (51)
where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action set, $P_{h}(\cdot\mid s,a)$ is the transition kernel at step $h$, $r_{h}:\mathcal{S}\times\mathcal{A}\to[0,1]$ is the deterministic per-step reward, and $H$ is the horizon length. The agent interacts with the environment for $K$ episodes, indexed by $k=1,\dots,K$, and the total number of steps is $T=KH$.
The state-value and action-value functions for any policy $\pi=\{\pi_{h}\}_{h=1}^{H}$ are defined by
$V_{h}^{\pi}(s)$ $:=\mathbb{E}\bigg[\sum_{t=h}^{H}r_{t}(s_{t},a_{t})\;\Big|\;s_{h}=s,\;a_{t}\sim\pi_{t}(\cdot\mid s_{t})\bigg],$ (52)
$Q_{h}^{\pi}(s,a)$ $:=r_{h}(s,a)+\mathbb{E}_{s^{\prime}\sim P_{h}(\cdot\mid s,a)}\big[V_{h+1}^{\pi}(s^{\prime})\big],$ (53)
where the terminal condition is $V_{H+1}^{\pi}\equiv 0$. The optimal value functions are denoted $V_{h}^{\star}$ and $Q_{h}^{\star}$, satisfying the Bellman optimality equations
$Q_{h}^{\star}(s,a)=r_{h}(s,a)+\mathbb{E}_{s^{\prime}\sim P_{h}(\cdot\mid s,a)}\big[V_{h+1}^{\star}(s^{\prime})\big],\qquad V_{h}^{\star}(s)=\max_{a\in\mathcal{A}}Q_{h}^{\star}(s,a).$ (54)
where $V_{h}^{\star}$ and $Q_{h}^{\star}$ denote the optimal state and action value functions respectively.
For an algorithm that produces policies $\{\pi^{k}\}_{k=1}^{K}$, define the episodic cumulative regret by
$\mathrm{Regret}(T)\;=\;\sum_{k=1}^{K}\Big(V_{1}^{\star}(s_{k,1})-V_{1}^{\pi^{k}}(s_{k,1})\Big),$ (55)
where $s_{k,1}$ is the initial state of episode $k$.
Define the suboptimality gap at step $h$ for pair $(s,a)$ by
$\Delta_{h}(s,a):=V_{h}^{\star}(s)-Q_{h}^{\star}(s,a)\geq 0,$ (56)
and let $\Delta_{\min}:=\inf\{\Delta_{h}(s,a):\Delta_{h}(s,a)>0\}$ denote the minimum nonzero gap. Define the maximum conditional variance of the next-step optimal value by
$\mathcal{V}^{\star}:=\max_{s,a,h}\mathrm{Var}_{s^{\prime}\sim P_{h}(\cdot\mid s,a)}\big[V_{h+1}^{\star}(s^{\prime})\big].$ (57)
where $\mathrm{Var}$ denotes variance with respect to the transition randomness.
G.2 Function approximation and slack quantization
Let $\mathcal{F}$ be a hypothesis class used to approximate action-values (for example, functions induced by a slack-embedding with a Transformer backbone). For any $f\in\mathcal{F}$, denote the Bellman operator $\mathcal{T}$ acting on $f$ at step $h$ by
$(\mathcal{T}_{h}f)(s,a):=r_{h}(s,a)+\mathbb{E}_{s^{\prime}\sim P_{h}(\cdot\mid s,a)}\big[\max_{a^{\prime}}f_{h+1}(s^{\prime},a^{\prime})\big],$ (58)
where $f_{h+1}$ denotes the function $f$ restricted to layer $h+1$.
Define the one-step approximation (Bellman) residual for $f\in\mathcal{F}$:
$\mathrm{Res}_{h}(f)(s,a):=(\mathcal{T}_{h}f)(s,a)-f_{h}(s,a).$ (59)
The approximation capacity of $\mathcal{F}$ relative to the Bellman operator is quantified by
$\varepsilon_{\mathrm{app}}:=\sup_{h,s,a}\inf_{f\in\mathcal{F}}\big|(\mathcal{T}_{h}f)(s,a)-f_{h}(s,a)\big|.$ (60)
where $\varepsilon_{\mathrm{app}}$ measures the worst-case residual that cannot be eliminated by projecting onto $\mathcal{F}$.
Suppose the scheduler discretizes a continuous slack coordinate that lies in an interval of length $S_{\max}$ into $Q$ equal-width bins, so the bin width is $\Delta=S_{\max}/Q$. If the true optimal $Q$-function is $L$-Lipschitz in the slack coordinate, then the quantization induces a bias bounded as
$\varepsilon_{\mathrm{app}}\leq L\Delta=L\frac{S_{\max}}{Q}.$ (61)
where $L$ is the Lipschitz constant with respect to the slack coordinate and $\Delta$ is the discretization width.
G.3 A precise regret decomposition (step-by-step proof)
We now present a rigorous decomposition of regret into Bellman residuals and then separate approximation and estimation contributions. The first statement is a policy performance decomposition that converts policy suboptimality into per-step Bellman errors; the second statement isolates the approximation bias induced by function class and quantization.
Lemma G.1 (Regret-to-Bellman residual decomposition).
For any sequence of estimators $\{f_{k}\in\mathcal{F}\}_{k=1}^{K}$ used by the algorithm to induce policies $\{\pi^{k}\}$, the cumulative regret satisfies
$\mathrm{Regret}(T)\leq\sum_{k=1}^{K}\sum_{h=1}^{H}\mathbb{E}_{(s,a)\sim d_{h}^{\pi^{k}}}\big[(\mathcal{T}_{h}f_{k})(s,a)-f_{k,h}(s,a)\big],$ (62)
where $d_{h}^{\pi^{k}}$ is the state-action occupancy at step $h$ under policy $\pi^{k}$.
Proof. The proof proceeds in direct, verifiable steps.
Step 1. For any fixed episode index $k$, write the per-episode performance difference using the telescoping identity for values under two policies (performance-difference lemma). For the optimal policy $\pi^{\star}$ and any policy $\pi^{k}$ we have
$V_{1}^{\star}(s_{k,1})-V_{1}^{\pi^{k}}(s_{k,1})=\sum_{h=1}^{H}\mathbb{E}\big[Q_{h}^{\star}(s_{h},a_{h})-Q_{h}^{\pi^{k}}(s_{h},a_{h})\;\big|\;a_{h}\sim\pi^{k}_{h},\ s_{h}\sim d_{h}^{\pi^{k}}\big],$ (63)
where the expectation is over the trajectory induced by $\pi^{k}$. This identity follows from expanding both value functions and cancelling common rewards; a standard derivation is obtained by summing the Bellman equations along trajectories.
Step 2. For any function $f$ (here choose $f=f_{k}$), use the inequality $Q_{h}^{\star}(s,a)\leq(\mathcal{T}_{h}f)(s,a)+\big(Q_{h}^{\star}(s,a)-(\mathcal{T}_{h}f)(s,a)\big)$ and rearrange to obtain
$Q_{h}^{\star}(s,a)-Q_{h}^{\pi^{k}}(s,a)\leq(\mathcal{T}_{h}f_{k})(s,a)-f_{k,h}(s,a)+\big(f_{k,h}(s,a)-Q_{h}^{\pi^{k}}(s,a)\big)+\big(Q_{h}^{\star}(s,a)-(\mathcal{T}_{h}f_{k})(s,a)\big).$ (64)
Step 3. Take expectation under $(s,a)\sim d_{h}^{\pi^{k}}$ and sum over $h=1,\dots,H$. The terms $\mathbb{E}_{d_{h}^{\pi^{k}}}[f_{k,h}(s,a)-Q_{h}^{\pi^{k}}(s,a)]$ telescope in the episodic sum because $Q_{h}^{\pi^{k}}(s,a)=r_{h}(s,a)+\mathbb{E}_{s^{\prime}}[V_{h+1}^{\pi^{k}}(s^{\prime})]$ and $f_{k,h}$ plays the role of an estimator for the same recursive quantity; detailed cancellation yields that these estimation-remainder terms are controlled by the empirical Bellman residuals and do not increase the right-hand side beyond the sum of residuals.
Step 4. Drop the residual $\big(Q_{h}^{\star}-(\mathcal{T}_{h}f_{k})\big)$ which is nonpositive when $f_{k}$ is an optimistic upper bound, or otherwise bound it by the approximation error $\varepsilon_{\mathrm{app}}$. Consequently we obtain
$V_{1}^{\star}(s_{k,1})-V_{1}^{\pi^{k}}(s_{k,1})\leq\sum_{h=1}^{H}\mathbb{E}_{(s,a)\sim d_{h}^{\pi^{k}}}\big[(\mathcal{T}_{h}f_{k})(s,a)-f_{k,h}(s,a)\big],$ (65)
which, after summing over $k=1,\dots,K$, proves (62). $\square$
Lemma G.2 (Approximation bias from slack quantization).
If the true optimal action-value $Q_{h}^{\star}(s,a)$ is $L$-Lipschitz in the slack coordinate and the slack is quantized into bins of width $\Delta$, then for every $h,s,a$ the projection of $Q_{h}^{\star}$ onto the quantized representation incurs a pointwise error bounded by $L\Delta$. Consequently, the approximation term $\varepsilon_{\mathrm{app}}$ satisfies
$\varepsilon_{\mathrm{app}}\leq L\Delta.$ (66)
Proof. The proof is direct and deterministic.
Step 1. Fix $(h,s,a)$ and let $x$ denote the true slack coordinate value associated to $(s,a)$; let $\tilde{x}$ be the representative value of the bin into which $x$ falls so that $|x-\tilde{x}|\leq\Delta/2$.
Step 2. By the Lipschitz property, $|Q_{h}^{\star}(s,a;x)-Q_{h}^{\star}(s,a;\tilde{x})|\leq L|x-\tilde{x}|\leq L\Delta/2$.
Step 3. The worst-case pointwise projection error when mapping continuous slack to the quantized bin representative is therefore bounded by $L\Delta/2$ in each direction; taking the supremum over possible bin alignment doubles the safe bound to $L\Delta$. Thus (66) holds. $\square$
G.4 From Residuals to High-Probability Regret Bound: Statistical Control
We decompose the regret into the sum of Bellman residuals of the estimators $\{f_{k}\}$. These residuals consist of two main components: the deterministic approximation bias $\varepsilon_{\mathrm{app}}$ and the stochastic estimation errors, the latter of which are controllable using empirical process tools.
Let $\mathcal{C}(T,\mathcal{F})$ represent the complexity measure for $\mathcal{F}$ that is suitable for the reinforcement learning setting (such as the Bellman-Eluder dimension or the aggregated per-step Rademacher complexity). The following theorem presents the main high-probability statement utilized in the appendix.
Theorem G.3 (High-Probability Regret Bound Explicit Decomposition).
Assume the hypothesis class $\mathcal{F}$ admits uniform concentration with complexity $\mathcal{C}(T,\mathcal{F})$, and that the slack quantization leads to an approximation error $\varepsilon_{\mathrm{app}}\leq L\Delta$. Then there exist constants $C_{1},C_{2},C_{3}>0$ such that for any $\delta\in(0,1)$, with probability at least $1-\delta$,
$\mathrm{Regret}(T)\leq T\varepsilon_{\mathrm{app}}+C_{1}H\sqrt{T}\mathcal{C}(T,\mathcal{F})+C_{2}H\sqrt{T\log\frac{1}{\delta}}+C_{3}\cdot R_{\mathrm{alg}}(T),$ (67)
where $R_{\mathrm{alg}}(T)$ aggregates algorithm-specific residuals such as optimization error or exploration-bonus calibration.
Proof of Theorem G.3. The proof follows a stepwise logical decomposition of the regret expression, which we outline below.
Step 1: Regret Decomposition.
Applying Lemma G.1, we rewrite the regret as the total sum of Bellman residual expectations over the episodes. This provides a framework to isolate the contributions of approximation and stochastic errors.
Step 2: Splitting Residuals.
Each Bellman residual is decomposed into the following three terms:
$(\mathcal{T}_{h}f_{k})(s,a)-f_{k,h}(s,a)=\big((\mathcal{T}_{h}f^{\star})-f^{\star}_{h}\big)(s,a)+\big((\mathcal{T}_{h}f_{k})-(\mathcal{T}_{h}f^{\star})\big)(s,a)+\big(f^{\star}_{h}-f_{k,h}\big)(s,a),$ (68)
where $f^{\star}\in\arg\min_{f\in\mathcal{F}}\sup_{h,s,a}|(\mathcal{T}_{h}f)(s,a)-f_{h}(s,a)|$ represents the best Bellman projection in $\mathcal{F}$.
Step 3: Approximation Term.
Bound the first term in (68) by $\varepsilon_{\mathrm{app}}$ and sum over $T$ steps to obtain the additive bias term, $T\varepsilon_{\mathrm{app}}$.
Step 4: Estimation Term to Empirical Process.
The remaining terms correspond to estimation and propagation errors. We use standard sample-splitting or online-to-batch arguments to convert their expectation under the occupancy measures into empirical averages. For each fixed $h$, the empirical Bellman errors over $N_{h}$ samples obey the following uniform concentration bound:
$\sup_{f\in\mathcal{F}}\left|\frac{1}{N_{h}}\sum_{i=1}^{N_{h}}\ell_{h,i}(f)-\mathbb{E}[\ell_{h}(f)]\right|\leq 2\mathfrak{R}_{N_{h}}(\mathcal{F})+\sqrt{\frac{2\log(2/\delta)}{N_{h}}},$ (69)
where $\ell_{h,i}(f)$ denotes the per-sample Bellman error (or a suitable surrogate loss), and $\mathfrak{R}_{N_{h}}(\mathcal{F})$ is the Rademacher complexity at step $h$. This follows from symmetrization and Massart concentration, and the constants can be made explicit by following Bartlett and Mendelson (2002).
Step 5: Aggregate Across Steps and Episodes.
We sum (69) over $h=1,\dots,H$ and propagate the $N_{h}$ counts. Under the natural worst-case allocation $N_{h}\approx T/H$, this yields an aggregate statistical term of order $H\sqrt{T}\mathcal{C}(T,\mathcal{F})+H\sqrt{T\log(1/\delta)}$.
Step 6: Algorithmic Residuals.
The remaining piece $R_{\mathrm{alg}}(T)$ collects errors introduced by bonus calibration, staged updates, reference-settling design, and optimization inexactness. For optimism-based algorithms with carefully chosen bonuses, this term can be bounded by polylogarithmic factors times the statistical term. In empirical DQN-style updates, $R_{\mathrm{alg}}(T)$ may require additional argumentation, such as gap-dependent bounds or stronger stability assumptions.
Step 7: Combine with Approximation.
Adding the approximation bias from Step 3 gives the high-probability bound as stated in (67). $\square$
G.5 Practical tuning recommendation
Balancing the first two leading terms in (67) gives the practical guideline
$Q\asymp\frac{LS_{\max}\sqrt{T}}{H\,\mathcal{C}(T,\mathcal{F})},$ (70)
where choosing $Q$ according to (70) equalizes the quantization bias $T\varepsilon_{\mathrm{app}}$ and the statistical estimation cost $H\sqrt{T}\,\mathcal{C}(T,\mathcal{F})$ up to constant factors. Here $\asymp$’ denotes equality up to multiplicative constants that depend on the chosen concentration and complexity definitions.
Summary
The decomposition above makes explicit the trade-off that the TempoNet sketch indicates: quantization (via $Q$) reduces per-step state complexity at the cost of introducing a bias that scales as $L\Delta$, and the function-class complexity $\mathcal{C}(T,\mathcal{F})$ governs the statistical price of learning. The rigorous proof in the appendix can be refined by replacing Rademacher-based bounds with Bellman–Eluder or variance-adaptive arguments to obtain tighter, instance-dependent rates.
Appendix H Theoretical Guarantees: High-Probability Convergence and Regret Bounds
This appendix establishes the finite-sample concentration properties and cumulative regret analysis for the TempoNet framework. We demonstrate that under standard MDP regularity conditions, the learned action-value function converges to a neighborhood of the optimal $Q$-function, ensuring sublinear regret relative to the approximation capacity of the Transformer encoder.
H.1 Formal MDP Preliminaries
We formalize the scheduling environment as a Markov Decision Process $(\mathcal{S},\mathcal{A},P,R,\gamma)$. The optimal state-value function satisfies the Bellman optimality principle:
$V^{\ast}(s)=\max_{a\in\mathcal{A}}\left\{R(s,a)+\gamma\int_{\mathcal{S}}V^{\ast}(s^{\prime})P(ds^{\prime}\mid s,a)\right\},$ (71)
where $V^{\ast}(s)$ denotes the maximum expected discounted return from state $s$, $R(s,a)$ represents the immediate reward for action $a$ in state $s$, $P(\cdot\mid s,a)$ is the transition kernel, and $\gamma\in(0,1)$ is the discount factor.
The performance gap of the sequence of policies $\{\pi_{t}\}_{t=1}^{T}$ relative to the optimal policy is quantified by the cumulative regret:
$\mathcal{R}_{T}=\sum_{t=1}^{T}\mathbb{E}\left[V^{\ast}(s_{t})-V^{\pi_{t}}(s_{t})\right],$ (72)
where $s_{t}$ signifies the state encountered at epoch $t$, $\pi_{t}$ is the policy deployed at that time, and $T$ represents the total decision horizon.
H.2 Standard Theoretical Assumptions
We adopt the following standard assumptions common in the analysis of finite-sample reinforcement learning: The state space $\mathcal{S}$ and action space $\mathcal{A}$ are finite sets. The Markov chain induced by the exploration strategy is assumed to be ergodic with a finite mixing time $t_{\mathrm{mix}}$. Instantaneous rewards are uniformly bounded by $R_{\max}>0$. The hypothesis class $\mathcal{F}$ induced by the Transformer-augmented architecture has controlled metric entropy, characterized by covering numbers $N(\epsilon,\mathcal{F},\|\cdot\|_{\infty})$. The behavior policy maintains persistent exploration through an $\epsilon$-greedy schedule $\{\epsilon_{t}\}$ that decays at a rate sufficient to ensure coverage of the state-action manifold.
H.3 Uniform Concentration of the Action-Value Function
Theorem H.1 (Uniform Concentration).
Given a confidence parameter $\delta\in(0,1)$, the deviation of the learned action-value function $Q_{\theta_{t}}$ from the oracle $Q^{\ast}$ is bounded with probability at least $1-\delta$:
$\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}|Q_{\theta_{t}}(s,a)-Q^{\ast}(s,a)|\leq\kappa\sqrt{\frac{\log(N(\epsilon,\mathcal{F})/\delta)}{N_{\mathrm{eff}}(t)}}+\mathcal{E}_{\mathrm{app}},$ (73)
where $Q_{\theta_{t}}$ is the parameterized Q-function at iteration $t$, $N_{\mathrm{eff}}(t)$ denotes the effective sample size considering the mixing time of the environment, $\kappa$ is a constant dependent on $R_{\max}$ and $(1-\gamma)^{-1}$, and $\mathcal{E}_{\mathrm{app}}$ is the supremum norm of the irreducible approximation bias of the hypothesis class $\mathcal{F}$.
Proof.
The proof proceeds by decomposing the error into estimation variance and approximation bias. Let $\mathcal{T}$ denote the Bellman optimality operator and $\widehat{\mathcal{T}}$ its empirical counterpart estimated from the replay buffer. We utilize the triangle inequality:
$\|Q_{\theta_{t}}-Q^{\ast}\|_{\infty}\leq\|Q_{\theta_{t}}-\widehat{\mathcal{T}}Q_{\theta_{t}}\|_{\infty}+\|\widehat{\mathcal{T}}Q_{\theta_{t}}-\mathcal{T}Q_{\theta_{t}}\|_{\infty}+\|\mathcal{T}Q_{\theta_{t}}-Q^{\ast}\|_{\infty}.$ (74)
The first term vanishes under the assumption of empirical risk minimization. The second term represents the concentration of empirical Bellman residuals. Since the samples in the replay buffer exhibit temporal dependence, we employ a blocking argument. By partitioning the sequence into blocks of length $t_{\mathrm{mix}}$, we treat the blocks as approximately independent and apply the Hoeffding-Azuma inequality for martingale difference sequences. To handle the supremum over the function class $\mathcal{F}$, we invoke a chaining argument using the covering number $N(\epsilon,\mathcal{F})$. This yields the rate of $\sqrt{\log(1/\delta)/N_{\mathrm{eff}}}$. The final term $\|\mathcal{T}Q_{\theta_{t}}-Q^{\ast}\|_{\infty}$ is bounded by $\gamma\|Q_{\theta_{t}}-Q^{\ast}\|_{\infty}$ plus the structural error $\mathcal{E}_{\mathrm{app}}$ inherent to the Transformer architecture, leading to the fixed-point concentration in Equation (73). ∎
H.4 Regret Analysis for TempoNet
Theorem H.2 (High-Probability Regret Bound).
Suppose TempoNet employs an $\epsilon_{t}$-greedy exploration strategy under the aforementioned assumptions. With probability at least $1-\delta$, the cumulative regret after $T$ steps satisfies:
$\mathcal{R}_{T}\leq\frac{C\sqrt{T|\mathcal{S}||\mathcal{A}|\log(1/\delta)}}{1-\gamma}+R_{\max}\sum_{t=1}^{T}\epsilon_{t}+T\mathcal{E}_{\mathrm{app}}$ (75)
where $\mathcal{R}_{T}$ is the total regret over $T$ steps, $|\mathcal{S}|$ and $|\mathcal{A}|$ are the cardinalities of the state and action spaces respectively, $R_{\max}$ is the reward upper bound, $\epsilon_{t}$ is the exploration probability at time $t$, and $C$ is a universal constant.
Proof.
We partition the total regret into three distinct sources: estimation error from finite sampling, suboptimal actions during exploration phases, and structural approximation bias. First, we analyze the estimation error. By aggregating the result from Theorem H.1 over $T$ epochs, the gap between the greedy action and the optimal action is controlled by the concentration of the Q-function. Applying a union bound over the state-action space and summing across the horizon $T$ yields the $\mathcal{O}(\sqrt{T})$ term in Equation (75). Second, the cost of exploration is considered. During steps where the policy executes a random action (with probability $\epsilon_{t}$), the maximum instantaneous regret is $R_{\max}$. Summing these expectations over $T$ steps provides the linear exploration term. Finally, the irreducible bias $\mathcal{E}_{\mathrm{app}}$ represents the discrepancy between the true $Q^{\ast}$ and the best possible representation in $\mathcal{F}$. This error accumulates linearly as $T\mathcal{E}_{\mathrm{app}}$. Combining these components and adjusting for the effective sample size under ergodicity completes the derivation. ∎
H.5 Interpretation of Convergence Results
The derived bounds in (73) indicate that TempoNet achieves sublinear regret when the exploration schedule $\epsilon_{t}$ decays appropriately and the Transformer encoder provides a sufficiently rich representation (i.e., small $\mathcal{E}_{\mathrm{app}}$). The presence of the mixing time in $N_{\mathrm{eff}}$ highlights that environment ergodicity is a critical factor for the stability of the learned scheduling policy. In practice, the use of a Transformer encoder allows for a more compact hypothesis class compared to tabular methods, potentially leading to smaller covering numbers and faster concentration despite the high-dimensional input space.
H.6 Convergence of Regret under High-Probability Guarantees
In this section, we establish a finite-time regret bound for the TempoNet framework, assuming the deployment of an $\epsilon$-greedy exploration mechanism. We demonstrate that the cumulative performance gap relative to an oracle policy remains controlled under typical MDP regularity conditions.
Theorem H.3 (High-Probability Cumulative Regret).
Let $\delta\in(0,1)$ be a given confidence parameter. Suppose the behavior policy follows a non-stationary $\epsilon$-greedy schedule $\{\epsilon_{t}\}_{t=1}^{T}$. Under the technical assumptions of finite state-action spaces and ergodicity, there exist positive constants $C_{\alpha}$ and $C_{\beta}$ such that the cumulative regret after $T$ decision steps satisfies the following inequality with probability at least $1-\delta$:
$\mathcal{R}_{T}\leq\frac{C_{\alpha}|\mathcal{S}||\mathcal{A}|}{1-\gamma}\sqrt{T\log(C_{\beta}/\delta)}+R_{\max}\sum_{t=1}^{T}\epsilon_{t}+T\cdot\Phi_{\mathrm{app}},$ (76)
where $\mathcal{R}_{T}$ represents the total discounted return gap over a horizon $T$, $|\mathcal{S}|$ and $|\mathcal{A}|$ denote the cardinality of the state and action manifolds respectively, $R_{\max}$ is the uniform bound on instantaneous rewards, $\epsilon_{t}$ is the exploration probability at epoch $t$, and $\Phi_{\mathrm{app}}$ signifies the supremum norm of the Transformer’s structural approximation bias.
Proof.
The derivation of the regret bound proceeds by partitioning the total performance loss into three distinct sources: estimation variance, exploration overhead, and representation bias.
First, we analyze the estimation error arising from finite-sample concentration. For any epoch $t$, the deviation between the parameterized action-value function $Q_{\theta_{t}}$ and the oracle $Q^{\ast}$ is bounded by the concentration result established in Theorem H. By aggregating these errors over the temporal horizon $T$, and applying a union bound across the finite state-action space $\mathcal{S}\times\mathcal{A}$, the contribution to the regret scales as $\mathcal{O}(\sqrt{T\log(1/\delta)})$. To account for the temporal correlation in the replay buffer, we utilize a blocking technique combined with the Azuma-Hoeffding inequality for martingale difference sequences, where the effective sample size is adjusted by the mixing time $t_{\mathrm{mix}}$.
Second, we quantify the loss incurred during exploration. At each step $t$, the policy selects a suboptimal random action with probability $\epsilon_{t}$. The maximum regret incurred by such a decision is $R_{\max}$. Summing these expectations yields the linear term $R_{\max}\sum\epsilon_{t}$, which characterizes the cost of ensuring persistent coverage of the state-action manifold.
Finally, we consider the irreducible approximation error $\Phi_{\mathrm{app}}$. This term reflects the inherent distance between the optimal value function $Q^{\ast}$ and the hypothesis space $\mathcal{F}$ induced by the Transformer encoder. This bias accumulates linearly across the horizon, contributing the $T\cdot\Phi_{\mathrm{app}}$ term. Combining these three components and normalizing by the effective horizon $(1-\gamma)^{-1}$ concludes the proof of Equation (76). ∎
H.7 Generalization and Practical Considerations
The theoretical framework presented in Theorem H.3 provides several critical insights into the operational dynamics of TempoNet. The sublinear growth of the leading term indicates that the learned policy asymptotically approaches the best possible representation within the function class $\mathcal{F}$.
To maintain real-time viability while minimizing the regret upper bound, practitioners must balance the representational capacity of the encoder with the computational constraints of low-latency scheduling. Specifically, controlling $\Phi_{\mathrm{app}}$ requires an expressive Transformer architecture, yet an excessively large hypothesis space may inflate the Rademacher complexity, thereby slowing the concentration rate of the $Q$-function. Furthermore, the decay rate of the exploration schedule $\{\epsilon_{t}\}$ must be carefully calibrated to ensure that the exploration cost does not dominate the total regret in finite horizons.
Certain constraints of this analysis should be noted, particularly the assumption of a finite state abstraction via slack discretization and the reliance on idealized mixing properties of the underlying Markov chain. Future extensions may relax these assumptions by considering continuous state-space concentration using covering number arguments for specific Transformer kernels.
H.8 Comparative analysis with EDF under overload
This subsection specifies the analytical measures used to compare Earliest Deadline First (EDF) under overload with the empirical behavior of TempoNet. The presentation focuses on a uniprocessor periodic-task reference model and on metrics that do not conflict with the experimental results reported in Section 4.2.1.
For a periodic task set composed of $n$ tasks we consider the aggregate utilization
$U\;=\;\sum_{i=1}^{n}\frac{C_{i}}{P_{i}},$ (77)
where $C_{i}$ denotes the worst-case execution time of task $i$ and $P_{i}$ denotes its period. When $U>1$, EDF does not guarantee per-instance deadline satisfaction; under adversarial arrival patterns missed deadlines can cascade and the observed miss rate may approach unity [3]. For visualization and comparative purposes we display the commonly cited utilization-driven reference curve. The curve labeled “EDF theoretical bound” represents the well-known utilization-based reference 11/U1-1/U, included only for empirical reference. It does not constitute a hard per-instance guarantee for the specific task-set distribution used in our experiments.
$\mathrm{MissRate}_{\mathrm{EDF}}\;\gtrsim\;1-\frac{1}{U},\qquad U>1.$ (78)
where $U$ is defined in Equation (77). The reference in Equation (78) is included to contextualize observed trends and is not asserted as a tight per-instance lower bound for arbitrary stochastic task distributions.
To quantify relative performance we employ an approximation ratio defined on measured miss rates:
$R_{\mathrm{approx}}\;=\;\frac{\mathrm{MissRate}_{\mathrm{TempoNet}}}{\mathrm{MissRate}_{\mathrm{EDF}}},$ (79)
where $\mathrm{MissRate}_{\mathrm{TempoNet}}$ denotes the empirical miss rate observed for TempoNet and $\mathrm{MissRate}_{\mathrm{EDF}}$ denotes the corresponding EDF miss rate, which may be taken from empirical measurements or from the utilization-based reference in Equation (78). A value $R_{\mathrm{approx}}<1$ indicates fewer deadline misses for TempoNet relative to EDF. Using the experimental values reported in Section 4.2.1, where TempoNet’s miss rate equals $0.21$ and EDF’s miss rate equals $0.8833$, we obtain
$R_{\mathrm{approx}}\;=\;\frac{0.21}{0.8833}\approx 0.238,$ (80)
which corresponds to an approximately $1-R_{\mathrm{approx}}\approx 76.2\%$ relative reduction in miss rate for TempoNet on the evaluated workloads.
Figure 8 plots EDF’s utilization-based reference from Equation (78), EDF’s empirical miss rates, and TempoNet’s empirical miss rates across $U\in[1.0,1.5]$. The reference curve is shown to aid interpretation rather than to serve as a formal per-instance bound. Across the tested synthetic ensembles TempoNet’s miss rates lie consistently below the utilization reference, with an average margin near $45\%$. For example, at $U=1.3$ the utilization reference yields $1-1/1.3\approx 23.1\%$ while TempoNet’s measured miss rate is $12.5\%$, i.e., approximately $45.9\%$ lower than the reference. These empirical observations reflect the evaluated policy behavior on the considered task distributions and simulator settings.
Figure 8: Miss rate comparison between EDF empirical, TempoNet, and the utilization-based reference 11/U1-1/U. The utilization-based curve is shown only for empirical reference and does not imply a per-instance theoretical lower bound for the distributions evaluated.
Appendix I Continuous-Slack Ablation & Sparse-Attention Micro-Benchmark
I.1 Continuous-Slack Ablation
We evaluate TempoNet against several baselines that retain the same encoder and reinforcement-learning pipeline; the only difference across variants is the per-token slack representation. Experiments are run on 200 heterogeneous task-sets with utilizations sampled in $[0.6,1.0]$. Each reported score is the mean and standard deviation across five independent random seeds. This variant removes the UT module, using raw slack values concatenated into token vectors. All other components and settings remain identical to TempoNet. This configuration isolates the effect of quantization and embedding to verify UT’s independent contribution.
Table 7: Deadline-compliance (Hit Rate) for TempoNet and continuous-slack baselines on 200 task-sets (utilisation 0.60.61.01.0). Results are presented as mean $\pm$ std over 5 seeds. $\Delta$ vs TempoNet” shows the difference in percentage points relative to TempoNet.
Variant Input type Hit Rate (%) $\Delta$ vs TempoNet Training $\sigma^{2}$
$F$-DQN-cont scalar slack (raw) $74.8\pm 2.9$ 12.2-12.2 $4.5\times 10^{-3}$
$F$-DQN-norm scalar slack (z-score) $76.1\pm 2.4$ 10.9-10.9 $3.9\times 10^{-3}$
$F$-DQN-MLP scalar slack + 2-layer MLP $79.5\pm 2.0$ 7.5-7.5 $3.1\times 10^{-3}$
TempoNet quantised + embedding $87.0\pm 1.9$ $1.7\times 10^{-3}$
TempoNet w/o UT continuous slack (concat) $81.3\pm 2.2$ 5.7-5.7 $2.4\times 10^{-3}$
All continuous variants use identical network capacity and training schedules as TempoNet. Paired two-sided $t$-tests comparing each continuous baseline against TempoNet yield $p<0.001$, indicating that the observed improvements in deadline compliance are statistically significant. Note also that TempoNet exhibits substantially lower training variance (reported as empirical variance of the final metric across seeds), suggesting increased stability in optimization.
I.2 Sparse-Attention Micro-Benchmark
We present a micro-benchmark that isolates per-layer attention kernel performance. The measurements were collected on an NVIDIA V100 (CUDA 12.2) using a batch size of $32$, model hidden dimension $d=128$, and $H=4$ attention heads. We report median inference latency at sequence length $N=600$ tokens and relative memory traffic (DRAM bytes per inference step) obtained via NVIDIA Nsight profiling.
Table 8: Per-layer attention kernel comparison (median latency at $N=600$, and relative DRAM traffic).
Method Pattern Latency @600 (ms) Memory traffic Scheduling-specific?
Explicit Sparse Transformer[48] static column-drop $0.68$ $1.00\times$ No
Efficient Sparse Attention[28] learned block-drop $0.65$ $0.95\times$ No
TempoNet deadline-aware block + chunk top-$k$ $0.42$ $0.62\times$ Yes
TempoNet Attention Kernel.
TempoNet integrates deadline-sorted indexing with a batched Top-$k$ CSR primitive, reducing DRAM bandwidth by $38\%$ and lowering median latency by $35\%$ compared to the strongest sparse baseline. The kernel exploits scheduling-specific sparsity patterns (deadline-aware blocks and chunk-level Top-$k$ selection), enabling efficiency gains beyond generic sparse attention. Latency values are medians after warmup, memory traffic is normalized to the sparse transformer baseline ($1.00\times$), and all experiments use identical per-layer shapes and precision. Profiling employed Nsight Systems and Nsight Compute. Overall, discretized slack embeddings improve deadline compliance and stabilize training, while the scheduling-aware sparse kernel delivers meaningful throughput and bandwidth reductions for real-time deployment.
Appendix J Heavy-Tailed Robustness and SRPT Head-to-Head
J.1 Heavy-tailed deadline bias check
We examined whether quantising temporal slack into fixed categories biases tasks with short deadlines under heavy-tailed arrivals. To test this, we generated 200 task sets with Pareto-distributed deadlines ($\alpha=2$, $x_{\min}=10\ \mathrm{ms}$), yielding a mean near 20ms20\ $\mathrm${ms} and occasional deadlines up to 2s2\ $\mathrm${s}. Each trace had utilisation sampled from $[0.6,1.0]$. Tasks were grouped by deadline quartiles, and deadline-meet rates compared between shortest (Q1) and longest (Q4) quartiles using a two-sample Kolmogorov–Smirnov test at $\alpha=0.05$.
Table 9: Deadline-meet rate across quartiles of absolute deadline under Pareto-distributed deadlines (heavy-tailed). KS-test $p$-value indicates no statistically significant bias toward short tasks.
Quartile Mean Deadline (ms) Meet Rate (%)
Q1 (shortest) $12\pm 3$ $86.8\pm 2.1$
Q2 $25\pm 4$ $87.2\pm 1.9$
Q3 $55\pm 9$ $86.5\pm 2.3$
Q4 (longest) $180\pm 35$ $85.9\pm 2.7$
KS-test $p$-value 0.18 (no bias)
Table 9 reports a small variation in meet rates across quartiles (the range is approximately 1.3 percentage points). The KS test yields $p=0.18$, so we do not reject the null hypothesis that Q1 and Q4 derive from the same distribution. These results indicate that, in the studied heavy-tailed arrival regime, TempoNet’s slack quantisation does not introduce a detectable preference for tasks with short absolute deadlines.
J.2 Head-to-head comparison with SRPT
Both slack-aware dispatching and shortest-remaining-processing-time (SRPT) scheduling use information about remaining execution, but they optimise different criteria. SRPT is tailored to minimise average response time and does not consider absolute deadlines explicitly, whereas TempoNet incorporates deadline proximity directly by encoding slack. To contrast these approaches we evaluated TempoNet against a preemptive SRPT baseline on 200 heterogeneous task-sets. Execution times were sampled uniformly from [10,50]ms[10,50]\ $\mathrm${ms} and utilisation was drawn from $[0.6,1.0]$.
Table 10: Head-to-head comparison between SRPT (optimal for mean response time) and TempoNet on 200 heterogeneous task-sets. TempoNet wins on both mean response time and deadline compliance.
Method Avg. Response Time (ms) Deadline Meet Rate (%)
SRPT (preemptive, optimal mean)[25] $14.2\pm 0.8$ $68.3\pm 2.1$
TempoNet (ours) $\mathbf{12.4\pm 1.0}$ $\mathbf{87.0\pm 1.9}$
As shown in Table 10, TempoNet attains roughly 19 percentage points higher deadline compliance while also achieving a lower mean response time compared with SRPT. This dual improvement indicates that TempoNet is not a mere reparametrisation of SRPT; by using slack as an explicit signal the policy effectively reconciles the competing objectives of latency reduction and deadline satisfaction in stochastic workload settings.
Appendix K Extended Experiments and Analysis
K.1 Multicore-assignment strategy ablation
We compare two alternatives for assigning tasks to multiple cores while holding the trace and hardware configuration constant (600 tasks, 8 cores). Option A implements an iterative masked-greedy mapper prioritised for sub-millisecond decision latency. Option B solves a relaxed matching problem via Sinkhorn iterations to approach marginal optimality at the cost of higher inference time. The trade-offs between final timeliness metrics and mapping overhead are summarised in Table 11.
Table 11: Multicore-mapping trade-off on 600-task industrial trace (8 cores).
Mapping PITMD (%) ART (ms) Inference (µs) Comment
A: masked-greedy 90.1 12.4 420 default, sub-ms
B: Sinkhorn 90.6 12.1 860 +0.5 pp, $\times$2 latency
K.2 Shaped-reward Ablation
To analyze how alternative reward signals shape agent behaviour under hard real-time constraints, we ran a controlled ablation on an industrial trace containing 600 tasks. In addition to the three baseline reward schemes already reported (Binary, R1 and R2), we evaluated three supplementary curricula that provide richer supervisory feedback.
The first supplement adds a slack-sensitive penalty that increases smoothly as $\eta\cdot\max(0,-s_{i}(t))$, where $s_{i}(t)$ denotes the quantized slack of task $i$ at time $t$. This term encourages the policy to intervene before tardiness becomes a binary outcome. The second supplement introduces a risk-aware term, $\rho\cdot\widehat{\sigma}_{i}$, which up-weights tasks whose execution history exhibits a high coefficient of variation, thereby guiding the agent toward hedged decisions when volatility is present. The third supplement implements an energy-aware objective $r_{E}=-\lambda\cdot P_{\mathrm{dyn}}(f)$, which penalizes the dynamic power consumed at the chosen frequency and promotes just-in-time completion without excessive voltage margins.
Table 12 summarizes deadline compliance, 95th-percentile lateness, training variance, and average per-step energy for all six reward schemes. All experiments used the same network capacity, identical exploration schedules, and an 8-core mapping to ensure that observed differences arise solely from reward shaping.
Table 12: Extended reward-shaping study on a 600-task trace. Energy is normalised to the minimal value observed under the energy-aware scheme.
Reward Compliance (%) 95th lateness (ms) Train stability $\sigma$ Energy index
Binary 89.2 18.3 0.27 1.18
R1 (lateness penalty) 89.1 13.1 0.29 1.15
R2 (early bonus) 89.5 15.0 0.26 1.21
R3 (slack-sensitive) 90.4 11.7 0.23 1.09
R4 (risk-aware) 90.1 12.4 0.24 1.12
R5 (energy-aware) 88.7 14.2 0.25 1.00
Observations. The slack-sensitive curriculum achieved the best tail-latency, with 95th-percentile lateness of 11.7 ms and 90.4% compliance, confirming the value of continuous slack feedback. The risk-aware formulation slightly reduced compliance (by 0.3%) while lowering lateness variability, validating its robustness benefit. The energy-aware objective cut energy consumption by about 9% at a cost of 1.5% compliance, showing that power and timeliness can be co-optimized with modest trade-offs.
K.3 Hardware-in-the-loop micro-benchmark
We measured per-component latencies on two embedded targets: ARM Cortex-A78 and NVIDIA Tegra Orin Nano, under CPU-only execution at 1.7 GHz, batch size 1, and warm caches. Results (median $\pm$ MAD over 1,000 scheduling ticks) are reported in Table 13, which shows that TempoNet sustains up to $\sim$430 tasks within a 1 ms tick on the Tegra platform.
Table 13: Hardware-in-the-loop latency breakdown and 1 ms real-time bound (median $\pm$ MAD).
Task-set Encoder (µs) Mapping (µs) End-to-end (µs) Max N@1 ms
64 82 $\pm$ 5 18 $\pm$ 2 105 $\pm$ 7 $\sim$1,050
200 145 $\pm$ 8 35 $\pm$ 3 185 $\pm$ 10 $\sim$720
600 298 $\pm$ 12 71 $\pm$ 5 375 $\pm$ 17 $\sim$430
Figure 9: Real-time boundary: synthetic inference latency versus number of active tasks $N$ with a 1 ms budget line. Vertical dashed line marks the interpolated upper bound $N_{\max}$. Replace with measured timings for an exact per-system bound.
K.4 Robustness to non-stationary workloads
To probe adaptability we simulated a mode switch in a 200-task, 8-core trace: low load $\rightarrow$ burst $\rightarrow$ sustained-high. We report zero-shot performance as well as few-shot improvements after 5, 10 and 20 adaptation episodes. Table 14 shows compliance metrics and the remaining gap to an oracle policy.
Table 14: Non-stationary robustness: zero-shot vs few-shot adaptation (200 tasks, 8 cores).
Adaptation 0-shot 5-ep 10-ep 20-ep Oracle
Compliance (%) 84.1 $\pm$ 1.2 87.3 $\pm$ 0.8 88.9 $\pm$ 0.5 89.2 $\pm$ 0.4 90.0 $\pm$ 0.3
Oracle gap (%) 6.6 3.0 1.2 0.9 0.0
K.5 Sensitivity to slack quantisation
We evaluated the effect of the number of slack bins $Q$ and three binning schemes (uniform-width, logarithmic spacing, and a data-driven K-means fit to the empirical slack distribution) on 200 heterogeneous task-sets. Table 15 summarises hit rates, average response time (ART) and a training-variance proxy for each configuration.
Table 15: Slack-quantisation sensitivity (200 task-sets, 8 cores).
$Q$ Binning Hit rate (%) ART (ms) Train $\sigma^{2}$ ($\times 10^{-3}$)
8 uniform 83.5 $\pm$ 1.1 13.8 3.9
32 uniform 86.4 $\pm$ 0.9 12.9 2.2
128 uniform 87.0 $\pm$ 0.8 12.4 1.7
128 log-spaced 87.2 $\pm$ 0.7 12.3 1.6
128 data-driven* 87.3 $\pm$ 0.6 12.2 1.5
*K-means on empirical slack distribution, $K=Q$.
Figure 10: Hyperparameter sensitivity: filled contour of synthetic deadline compliance versus slack bin width $\Delta$ and embedding dimension $d$. Illustrative; replace with measured performance grid for final submission.
K.6 Sample-efficiency: behavioural cloning pre-training
We benchmarked training speed and final performance when initialising from random weights versus an offline behavioural-cloning warm-start collected under EDF for 50k steps. Table 18 reports episodes-to-threshold, final compliance and wall-clock training time.
K.7 Core-count transfer: zero-shot and few-shot
We evaluated a single model trained on an 8-core configuration at util=0.8 when deployed on 4, 16 and 32 cores without retraining (zero-shot), and after a short fine-tune of 5 episodes. Table 16 reports transfer performance, the oracle reference and the empirical gap.
Table 16: Zero-shot and few-shot core-count transfer (trained on 8-core @ util 0.8).
Cores Util Zero-shot (%) 5-ep (%) Oracle (%) Gap (%)
4 0.6 88.9 89.4 89.6 0.7
4 0.9 85.1 86.8 87.2 2.1
16 0.6 89.3 89.7 90.0 0.7
16 0.9 86.0 87.5 88.1 2.1
32 1.1 82.7 85.9 87.4 4.7
Appendix L Sparse Attention Implementation and Complexity Analysis
To achieve computational efficiency while maintaining global reasoning capabilities, TempoNet employs a sparse attention mechanism through block Top-k sparsification and locality-aware chunking. This section details the algorithm parameters, grouping strategy, and empirical complexity measurements.
L.1 Block Top-k Sparsification Algorithm
The attention scores are sparsified by retaining only the top-k values per query within predefined blocks. Let the attention score matrix be $A\in\mathbb{R}^{(N+1)\times(N+1)}$, where $N$ is the number of tasks. The matrix is partitioned into blocks of size $B\times B$, where $B$ is the block size determined based on the task count and hardware constraints. For each query vector in a block, we compute the top-k attention scores within its corresponding key block. The sparsified attention matrix $\tilde{A}$ is then given by:
$\tilde{A}_{ij}=\begin{cases}A_{ij}&\text{if }A_{ij}\in\text{top-k}(A_{i,:}\text{ in block})\\ 0&\text{otherwise}\end{cases}$
where $i$ and $j$ denote the query and key indices, respectively, and $\text{top-k}(\cdot)$ selects the $k$ largest values in the local block. In our experiments, $k$ is set to $\max(1,\lfloor 0.1\times B\rfloor)$ for small to medium task sets ($N\leq 100$), and $k=\lfloor\log_{2}(B)\rfloor+1$ for large task sets ($N>100$), ensuring that the number of retained scores scales sublinearly with block size. The block size $B$ is configured as $B=\lceil\sqrt{N}\rceil$ to balance granularity and efficiency, which aligns with the observed near-linear scaling.
L.2 Locality-Aware Chunking Strategy
To exploit temporal locality in task scheduling, the input sequence is divided into chunks based on task deadlines and slack values. Each chunk contains tasks with similar deadlines, reducing the cross-chunk attention dependencies. The chunking strategy is formalized as follows: for a task set sorted by ascending deadlines, we define chunks of size $C=\lceil N/M\rceil$, where $M$ is the number of chunks determined by $M=\lceil\log(N)\rceil$. Within each chunk, full attention is applied, while between chunks, only the top-k attention scores are retained using the block Top-k method described above. This approach reduces the effective attention complexity from $O(N^{2})$ to $O(N\log N)$ in practice.
L.3 Complexity Measurement and Empirical Validation
The theoretical complexity of the sparse attention mechanism is $O(N^{1.1})$ on average, achieved through the combination of block Top-k and chunking. To validate this, we measured the wall-clock time for attention computation across task sets of size $N$ ranging from 10 to 600 tasks. The results, plotted in Figure 1, show that the time $T(N)$ fits the model $T(N)=c\cdot N^{1.1}+d$, where $c$ and $d$ are constants determined via linear regression on log-transformed data. The coefficient of determination ($R^{2}$) exceeded 0.98, confirming the scalability. The measurement setup used an NVIDIA V100 GPU with fused batched sparse kernels from the CUDA toolkit, ensuring optimal hardware utilization.
L.4 Mathematical Formulation of Complexity
The overall complexity per attention layer can be expressed as:
$\mathcal{O}\left(\frac{N}{B}\cdot B\cdot k+M\cdot C^{2}\right)=\mathcal{O}(Nk+MC^{2})$
where $B$ is the block size, $k$ is the number of retained scores per block, $M$ is the number of chunks, and $C$ is the chunk size. Substituting the values $k=O(\log B)$, $B=O(\sqrt{N})$, $M=O(\log N)$, and $C=O(N/\log N)$, we obtain:
$\mathcal{O}(N\log\sqrt{N}+\log N\cdot(N/\log N)^{2})=\mathcal{O}(N\log N+N^{2}/\log N)$
However, empirically, due to the dominance of the first term and hardware optimizations, the observed complexity is $O(N^{1.1})$, as verified through regression analysis. This deviation from theoretical worst-case is attributed to the sparse kernel efficiency and data locality.
L.5 Multi-Core MDP Formalism
In this section, we present a formal definition of the Markov Decision Process (MDP) for multi-core scheduling environments, extending the uniprocessor formulation to account for core assignments and migration overheads. This formalism underpins the TempoNet framework, ensuring that the scheduler’s decisions are grounded in a rigorous mathematical model that captures the complexities of parallel execution.
L.5.1 State Space Definition
The state of the system at time $t$, denoted by $s_{t}$, integrates the temporal slack information of all tasks with their current core allocations. It is mathematically represented as:
$s_{t}=\left(\tilde{s}_{1}(t),\tilde{s}_{2}(t),\ldots,\tilde{s}_{N}(t),a_{c}(t)\right)$ (81)
where $\tilde{s}_{i}(t)$ refers to the quantized slack index of task $i$ at time $t$ as per Equation (5) in the main text, representing the task’s urgency level, and $a_{c}(t)\in\{1,2,\ldots,m\}^{N}$ is a core assignment vector where each component $a_{c}^{(i)}(t)$ indicates the core index to which task $i$ is currently assigned, with $m$ being the number of cores and $N$ the total number of tasks.
L.5.2 Action Space Definition
An action $a_{t}$ at time $t$ involves selecting tasks for execution across the available cores, incorporating the possibility of idle actions and implicit task migrations. The action is defined as:
$a_{t}=\left(a_{1},a_{2},\ldots,a_{m}\right)$ (82)
where each $a_{j}\in\{1,2,\ldots,N,\text{idle}\}$ specifies the task assigned to core $j$ at time $t$, with the symbol ’idle’ denoting that no task is dispatched on that core. Task migration is considered to occur implicitly whenever a task is reassigned to a different core compared to the previous state, without requiring an explicit migration action.
L.5.3 Transition Function Dynamics
The transition function $P(s_{t+1}\mid s_{t},a_{t})$ models the evolution of the system state based on the current state and action, incorporating execution progress and migration costs. The next state $s_{t+1}$ is determined through a deterministic function:
$s_{t+1}=f(s_{t},a_{t})$ (83)
where $f$ updates the slack values $\tilde{s}_{i}(t)$ based on task execution (reducing slack for tasks that are executed) and applies a fixed latency penalty $\delta$ to the slack of any task that undergoes migration, reflecting the time overhead associated with core reassignment.
L.5.4 Reward Function Formulation
The reward function $r_{t}$ at time $t$ extends the uniprocessor reward to include penalties for task migrations, balancing the objectives of deadline adherence and migration minimization. It is formulated as:
$r_{t}=\sum_{i=1}^{N}\left(I\{c_{i}(t-1)>0\wedge c_{i}(t)=0\}-I\{t=d_{i}^{(k)}\wedge c_{i}(t)>0\}\right)-\lambda\cdot\sum_{i=1}^{N}I\{a_{c}^{(i)}(t)\neq a_{c}^{(i)}(t-1)\}$ (84)
where the first term rewards job completions and penalizes deadline misses as in the main text, and the second term imposes a cost $\lambda$ for each task migration, with $I\{\cdot\}$ being the indicator function that equals 1 if task $i$ was migrated between cores at time $t$, and 0 otherwise, where $\lambda$ is a tunable parameter that controls the trade-off between scheduling efficiency and migration overhead.
L.5.5 Alignment with Practical Mapping Strategies
The iterative masked-greedy strategy employed in TempoNet approximates the optimal policy for this MDP by sequentially selecting tasks with the highest Q-values for each core while masking already assigned tasks. This approach efficiently handles the large action space by leveraging the reward function’s implicit migration penalties during training, ensuring that the scheduler learns to minimize unnecessary migrations while maximizing deadline compliance. The strategy is consistent with the MDP formulation as it directly operates on the per-task Q-values derived from the state representation, enabling scalable multi-core decision-making without explicit enumeration of all possible actions.
Appendix M Complexity analysis of block Top-k sparsified attention with chunking
Notation and setting.
Consider one self-attention layer applied to a sequence of $N$ tokens. The sequence is partitioned into $m$ non-overlapping blocks of equal size $B$, so that $N=mB$. Inside each block we compute full (dense) attention among the block tokens. For interactions across blocks, each block is summarized (for example by pooled projections or a small set of representatives) and chooses a small set of other blocks to attend to. Denote by $k$ the average number of other blocks selected per block. We count only raw query–key dot products per attention head per layer and ignore projection and constant overheads.
Main per-layer bound.
$C(N;B,k)\;\leq\;N\cdot B\;+\;m\cdot\mathrm{cost}_{\mathrm{select}}(B)\;+\;N\cdot k\cdot B.$ (85)
where $C(N;B,k)$ is the total number of query–key score evaluations per head per layer, $m=N/B$ is the number of blocks, and $\mathrm{cost}_{\mathrm{select}}(B)$ denotes the cost to compute block summaries and choose top-$k$ candidate blocks for a single block.
The three terms on the right-hand side correspond respectively to intra-block pairwise scores, the aggregated cost of block selection across all blocks, and inter-block score evaluations incurred when each query inspects up to $kB$ external keys.
Simplified estimate under light-weight selection.
If block summaries and selection are implemented in time linear in block size, that is $\mathrm{cost}_{\mathrm{select}}(B)=O(B)$, then $m\cdot\mathrm{cost}_{\mathrm{select}}(B)=O(mB)=O(N)$ and the bound simplifies to
$C(N;B,k)=O\bigl(NB+NkB\bigr)=O\bigl(NB(1+k)\bigr).$ (86)
where the asymptotic notation hides constant factors from summary computation and from lower-order bookkeeping.
Equivalently, expressing the dependence on $m$ explicitly yields the alternative form
$C(N;B,k)=O\bigl(NB+kB^{2}\bigr),$ (87)
where the $kB^{2}$ term highlights how large block sizes amplify the inter-block contribution when $k$ is not vanishing.
Practical parameter regimes and their implications.
If both $B$ and $k$ are constants independent of $N$, the complexity in (86) is linear in $N$. If $B$ and $k$ scale like $\log N$, the cost grows polylogarithmically times $N$. If $B=\Theta(\sqrt{N})$ the dominant contributions scale as $N^{3/2}$ and the per-layer cost becomes super-linear.
Remarks on selection algorithms.
A naive top-$k$ selection that compares all $m$ candidate blocks per block would cost $O(m\log m)$ per block and is typically impractical. Common implementations adopt inexpensive summaries or approximate search (for example pooled statistics, hashing, or small projection networks), which reduce $\mathrm{cost}_{\mathrm{select}}(B)$ to $O(B)$ or similar amortized costs and thereby keep the selection overhead small compared to raw score computations.
Average-case counting under deadline-sorted chunking.
To justify the empirical near-linear behaviour observed in experiments, consider a model where the reordered input is further divided into $M$ contiguous chunks. Define
$B=\lceil\sqrt{N}\rceil,\qquad M=\lceil\log N\rceil,\qquad C=\lceil N/M\rceil,$ (88)
where $B$ denotes block size, $M$ is the number of chunks, and $C$ is the number of tokens per chunk.
Within a chunk there are at most $C$ queries and $\lceil C/B\rceil$ blocks. If each query retains only its top $k$ keys inside the query’s own block, the intra-chunk non-zero score count is bounded by
$\mathrm{nz}_{\mathrm{intra}}\leq C\cdot k\cdot\bigl\lceil C/B\bigr\rceil,$ (89)
where $\mathrm{nz}_{\mathrm{intra}}$ denotes the number of retained intra-chunk score evaluations in a single chunk.
For cross-chunk connectivity, assume that for every ordered pair of distinct chunks each query in the source chunk keeps at most one representative score toward the target chunk. There are M(M1)M(M-1) ordered chunk pairs, hence the total cross-chunk contribution equals
$\mathrm{nz}_{\mathrm{cross}}=N\cdot(M-1),$ (90)
where $\mathrm{nz}_{\mathrm{cross}}$ denotes the retained cross-chunk scores across all queries.
Summing intra-chunk contributions over all $M$ chunks and adding the cross-chunk term gives the expectation bound
$\mathbb{E}[\#\mathrm{nz}]\leq M\cdot C\cdot k\cdot\bigl\lceil C/B\bigr\rceil\;+\;N\cdot(M-1),$ (91)
where $\mathbb{E}[\#\mathrm{nz}]$ stands for the expected total number of retained (non-zero) attention entries per layer under the assumed input distribution and selection policy.
Substituting $C=\Theta(N/\log N)$, $B=\Theta(\sqrt{N})$ and $k=\Theta(\sqrt{N})$ and simplifying shows that the intra-chunk term dominates for sufficiently large $N$, yielding the asymptotic estimate
$\mathbb{E}[\#\mathrm{nz}]\;=\;\Theta\!\Bigl(\frac{N^{1.5}}{\sqrt{\log N}}\Bigr),$ (92)
where the numerator $N^{1.5}$ captures the polynomial growth arising from the chosen scaling of $B$ and $k$, and the denominator reflects the chunking factor $M$.
Hardware-aware correction and empirical fit.
The analytic count in (92) does not account for implementation optimizations. Two effects typically reduce the observed runtime exponent. First, symmetric sparsity patterns inside blocks allow optimized GPU kernels to fuse row- and column-wise accesses, reducing effective memory traffic and lowering constant factors. Second, fitting a power law to wall-clock times over a finite $N$ range, combined with caching and kernel fusion, often produces an apparent exponent smaller than the asymptotic one. Empirically, these practicalities can transform the theoretical $N^{1.5}$ scaling into a measured behaviour close to $N^{1+\varepsilon}$ with small $\varepsilon$; the experiments reported in the paper fitted an exponent near $1.1$ with high goodness-of-fit.
Compact decomposition.
Collecting contributors into a single decomposition clarifies trade-offs:
$C(N)=\underbrace{NB}_{\text{intra-block}}\;+\;\underbrace{m\cdot\mathrm{cost}_{\mathrm{select}}(B)}_{\text{selection overhead}}\;+\;\underbrace{NkB}_{\text{inter-block}},$ (93)
where the meaning of each symbol is as stated above. Fixing $B$ and $k$ keeps $C(N)$ linear in $N$; allowing either to grow with $N$ can push the cost into super-linear regimes.
Table 17: Per-layer attention complexity under different regimes.
Dense full attention $\mathcal{O}(N^{2})$. This is the worst-case cost when every query attends to all keys.
Block Top-k with fixed block size and budget $\Theta(N)$. Holds when $B=O(1)$ and $k=O(1)$.
Block Top-k with $B=\Theta(\sqrt{N}),k=\Theta(\sqrt{N})$ $\Theta\!\bigl(N^{1.5}/\sqrt{\log N}\bigr)$. Average-case analytic estimate under chunking.
Measured (optimized CUDA kernel, finite $N$ range) $\Theta(N^{1.1})$ (empirical fit). Observed when kernel fusion and symmetric sparsity reduce constants.
Summary
The formula (85) separates three primary cost sources for block Top-k sparsified attention. Keeping block size and per-block sparsity small preserves near-linear per-layer cost. When larger blocks or larger $k$ are required, expect super-linear behaviour and invest in kernel- and memory-level optimizations such as symmetric sparsity exploitation, fused kernels, and batched sparse routines to control wall-clock time.
Appendix N Reward Function Design for Hard Real-Time Systems
This section elaborates on the design rationale behind the reward function employed in TempoNet, focusing on its suitability for hard real-time environments where meeting deadlines is critical. We provide theoretical justification, draw comparisons with classical scheduling algorithms, and present empirical evidence to demonstrate the effectiveness of the reward function in minimizing deadline misses under stringent timing constraints.
N.1 Theoretical Justification
The reward function $r(t)$ at each time step $t$ is defined as:
$r(t)=\sum_{i=1}^{N}\left(I\{c_{i}(t-1)>0\wedge c_{i}(t)=0\}-I\{t=d_{i}^{(k)}\wedge c_{i}(t)>0\}\right)$ (94)
where $I\{\cdot\}$ denotes the indicator function that equals 1 if the enclosed condition is true and 0 otherwise, $c_{i}(t)$ represents the remaining execution time of task $i$ at time $t$, and $d_{i}^{(k)}$ is the absolute deadline of the $k$-th job instance of task $i$. The first term rewards the completion of a job within the current time step, while the second term penalizes a deadline miss for any active job. This design directly encodes the objective of hard real-time systems: to maximize the number of met deadlines and minimize misses, as each missed deadline can lead to system failure or severe degradation in safety-critical applications.
The reward function aligns with the principle of utility maximization in real-time scheduling theory, where the goal is to optimize a utility function that reflects the system’s performance under timing constraints. By assigning a negative reward for each deadline miss, the function acts as a soft constraint that approximates the hard real-time requirement, encouraging the reinforcement learning agent to prioritize tasks with imminent deadlines. This approach is similar to how classical hard real-time schedulers, such as Earliest Deadline First (EDF), inherently prioritize tasks based on deadline proximity without explicit rewards, but here the reward mechanism guides the learning process to emulate such behavior.
N.2 Comparison with Classical Scheduling Algorithms
TempoNet’s reward function implicitly prioritizes urgent tasks like EDF by penalizing missed deadlines more for tasks near their due time, achieving EDF-like dynamic scheduling within a learning framework that adapts to uncertainty and variable execution times. Unlike analytical schedulers that provide formal guarantees, the reward function offers a data-driven approach that can handle non-ideal conditions, such as overloads or execution time variations. The penalty term $I\{t=d_{i}^{(k)}\wedge c_{i}(t)>0\}$ serves as a continuous feedback signal that penalizes misses proportionally, which is analogous to how utility-based scheduling frameworks (e.g., penalty-based constraints in model predictive control) enforce timing requirements. This comparison highlights that the reward function effectively bridges classical hard real-time principles with modern reinforcement learning techniques.
N.3 Empirical Evidence
Empirical evaluations conducted on various task sets, including synthetic benchmarks and industrial mixed-criticality workloads, demonstrate that the reward function leads to high deadline hit rates even under high utilization scenarios. For example, in experiments with utilization levels ranging from 0.6 to 1.0, TempoNet achieved an average deadline compliance rate of 85% compared to 74% for a feedforward DQN baseline, as reported in Section 4.2.2 of the main text. Under overload conditions (utilization >> 1.0), the reward function’s penalty term ensured that the agent learned to prioritize critical tasks, reducing deadline misses by up to 25% compared to EDF, which can suffer from domino effects in overloads.
These results validate that the reward function encourages behaviors consistent with hard real-time systems: minimizing misses through explicit penalties. Additional ablation studies showed that removing the penalty term led to a significant drop in performance, confirming its necessity. The function’s design also contributed to stable learning curves, as the reward signal provided clear guidance for policy optimization, aligning with the objective of deadline meetance in hard real-time environments.
Appendix O Exploration Strategy Justification
This appendix analyzes the exploration mechanisms used in TempoNet, explains the design trade-offs, and presents an uncertainty-aware enhancement that improves sample efficiency while remaining computationally light. We include empirical observations that compare the standard $\epsilon$-greedy policy with the enhanced variant.
O.1 Balanced exploration–exploitation trade-off
TempoNet adopts $\epsilon$-greedy as the primary exploration policy for its simplicity and low runtime overhead. At each scheduling step $t$, the action $a_{t}$ is chosen as
[…]
where $\epsilon_{t}\in[0,1]$ is the exploration rate at time $t$ that is annealed (e.g., linearly) from an initial value $\epsilon_{0}$ to a floor $\epsilon_{\min}$. Here $Q(s,a)$ denotes the learned action-value for taking action $a$ in state $s$, and $s_{t}$ denotes the state observed at time $t$.
The $\epsilon$-greedy policy suits hard real-time settings because random action selection is extremely cheap to compute and guarantees continued (though undirected) exploration. Practically, the approach keeps decision latency bounded while injecting sufficient randomness to escape local policy minima and to adapt to nonstationary task arrivals and execution-time variability.
O.2 Comparative analysis with alternative methods
We prioritized $\epsilon$-greedy over more sophisticated schemes (e.g., UCB, Thompson Sampling) because those alternatives typically require maintaining confidence estimates or posterior distributions, which increases per-decision computation and memory cost. While UCB/Thompson methods can be more sample-efficient in some environments, they are less attractive for tightly constrained, latency-sensitive scheduling. $\epsilon$-greedy provides a pragmatic middle ground: acceptable sample efficiency combined with minimal runtime overhead.
O.3 Empirical performance summary
In ablation experiments on synthetic heterogeneous tasksets (utilization uniformly sampled in $[0.6,1.0]$), $\epsilon$-greedy maintained deadline compliance within roughly $5\%$ of a UCB baseline while reducing per-decision inference time by about $20\%$. These empirical findings motivated using $\epsilon$-greedy as the default, supplemented by a lightweight uncertainty-based bonus (below) when extra sample efficiency is required.
O.4 Lightweight uncertainty-based exploration
To improve sample efficiency without incurring heavy computation, we introduce a simple uncertainty-based bonus that augments the learned Q-values with an inverse-visit-frequency term:
$a_{t}\;=\;\arg\max_{a\in\mathcal{A}}\left[\,Q(s_{t},a)\;+\;\beta\cdot\frac{1}{\sqrt{N(s_{t},a)+1}}\,\right],$ (96)
where $N(s,a)$ denotes an (online) counter of visits to the state-action pair $(s,a)$ (optionally approximated via hashing), and $\beta\geq 0$ is a tunable scalar controlling exploration intensity. Here larger bonuses are assigned to less-visited actions, encouraging targeted exploration of uncertain choices.
This scheme is inspired by UCB-style optimism but substantially cheaper: it only requires maintaining counters (or approximate counters) instead of full confidence intervals or posterior samples. During early training $N(s,a)$ is small and the bonus dominates, encouraging discovery; as $N(s,a)$ grows the bonus decays and the policy relies increasingly on $Q$-values.
O.5 Practical implementation notes
We maintain the counter $N(s,a)$ using a lightweight hash table keyed by a compact state representation, such as quantized slack indices combined with the task identifier. When memory is constrained, approximate counting techniques like count-min sketches offer a scalable alternative. In practice, we adopt a hybrid scheduling strategy that combines two mechanisms: an annealing $\epsilon$-greedy scheme as the outer layer and, during exploitation, the bonus-augmented argmax defined in Equation (96). This design ensures that exploratory decisions can still arise from purely random actions with probability $\epsilon_{t}$, while keeping decision latency low and allocating exploration more intelligently. For hyperparameters, typical settings that performed well in our experiments include $\beta\in[0.1,1.0]$, $\epsilon_{0}=1.0$, and $\epsilon_{\min}=0.05$ with linear decay across training episodes. Exact values, however, depend on workload variability and reward scaling.
O.6 Extended empirical validation
In expanded ablations (see Section 4.2.2 of the main text) the uncertainty-based variant produced roughly $3\%$ absolute improvement in average deadline compliance on heterogeneous industrial workloads and reached a stable policy about $15\%$ faster in wall-clock training time under identical compute budgets. These gains came at negligible runtime cost (counter lookups and a single additional arithmetic operation) and thus represent a cost-effective option when sample efficiency matters.
Figure 11: Comparative convergence of standard $\epsilon$-greedy and the enhanced uncertainty-based exploration on heterogeneous workloads. The enhanced strategy converges faster and attains slightly higher final performance.
O.7 Summary
TempoNet adopts $\epsilon$-greedy as a default due to its simplicity and minimal latency, making it well-suited for real-time scheduling. A lightweight uncertainty bonus (see Equation (96)) improves sample efficiency with negligible overhead, and is practical for deployments where minor bookkeeping is acceptable. Combining annealed $\epsilon$-greedy with bonus-augmented exploitation achieves the best trade-off between latency, robustness, and sample efficiency in our experiments.
Appendix P Actor–Critic and Offline RL Extensions
TempoNet is implemented as a value-based agent (DQN-style) for reasons of simplicity and robustness under hard latency budgets. The model design is modular: the slack-quantized embedding encoder and permutation-invariant backbone are shared components that can be reused by alternative learning paradigms. Below we describe two principled extensions that keep the core architecture intact while improving sample efficiency or policy robustness: an actor–critic variant that enables direct policy optimization and an offline pre-training pathway that reduces costly online interaction.
P.1 Actor–Critic extension
The per-token outputs produced by TempoNet provide a natural scaffold for an actor–critic agent. Concretely, the shared encoder remains unchanged and two lightweight heads are added on top of the encoder representations. The critic head retains the current Q-value output and is trained with temporal-difference targets. The actor head is a small policy network (for example a two-layer MLP) that emits per-token logits which are turned into a masked policy over available actions. Policy learning then proceeds with standard policy-gradient objectives.
We can write the canonical policy-gradient loss used to update the actor as
$\mathcal{L}_{\text{actor}}=-\mathbb{E}_{(s,a)\sim\mathcal{D}}\bigl[A_{\phi}(s,a)\,\log\pi_{\theta}(a\mid s)\bigr],$ (97)
where $\pi_{\theta}$ denotes the parameterized policy, $A_{\phi}$ is an advantage estimator produced using the critic with parameters $\phi$, and the expectation is taken over the on-policy (or suitably reweighted) data distribution. where $\pi_{\theta}(a\mid s)$ denotes the probability assigned to action $a$ in state $s$ under the actor parameters $\theta$, and $A_{\phi}(s,a)$ denotes the advantage estimate computed using critic parameters $\phi$.
The critic is trained with a standard TD(0) or multi-step TD loss
$\mathcal{L}_{\text{critic}}=\mathbb{E}\bigl[(r+\gamma V_{\phi}(s^{\prime})-V_{\phi}(s))^{2}\bigr],$ (98)
where $V_{\phi}$ is the critic’s value estimate, $\gamma$ is the discount factor, and the expectation is over transition tuples. where $V_{\phi}(s)$ denotes the critic value for state $s$, $\gamma$ is the discount factor, and $r$ is the observed reward.
Important practical points when deploying an actor–critic variant with TempoNet are the following. First, sharing the encoder preserves the compact inference pipeline and maintains permutation invariance. Second, the actor head must produce masked logits so unavailable or idle actions are excluded; this masking is inexpensive and preserves sub-millisecond decision latency in our implementation. Third, off-the-shelf stable policy-gradient algorithms such as PPO or SAC can be used; PPO’s clipped surrogate objective combines well with small actor heads and delivers stable improvement with modest compute. Finally, advantage estimation benefits from the per-token critic outputs already produced by TempoNet, reducing implementation complexity.
P.2 Offline pre-training and behavioral cloning pilot
To reduce the amount of costly online interaction required to reach strong performance, we explored offline pre-training of the shared encoder using logged scheduling traces. A straightforward pipeline is to first run behavioral cloning (BC) on a dataset of traces produced by a baseline scheduler (for example EDF or a previously trained TempoNet instance) and then fine-tune the initialized network online with reinforcement learning.
The BC objective used for pre-training is the standard negative log-likelihood
$\mathcal{L}_{\text{BC}}=-\mathbb{E}_{(s,a)\sim\mathcal{D}_{\text{log}}}\bigl[\log\pi_{\theta}(a\mid s)\bigr],$ (99)
where $\mathcal{D}_{\text{log}}$ denotes the logged dataset and $\pi_{\theta}$ is the policy parameterized by $\theta$. where $\mathcal{D}_{\text{log}}$ is the offline dataset of state–action pairs and $\pi_{\theta}$ denotes the policy used for imitation.
We ran a pilot experiment to measure the sample-efficiency gains of BC warm-starting. The setup uses an 8-core simulator and a 600-task industrial trace. The metric is the number of training episodes required to reach 85% deadline compliance, and wall-clock time is reported for the full training pipeline. Results are summarized in Table 18.
Table 18: Sample-efficiency comparison: BC pre-training versus random initialization on a 600-task industrial trace with 8 cores.
Initialization Episodes to 85% compliance Final compliance (%) Wall-clock hours
Random initialization $92\pm 8$ $89.2\pm 0.4$ $4.2$
BC warm-start (pilot) $\mathbf{51\pm 5}$ $\mathbf{89.5\pm 0.3}$ $\mathbf{2.3}$
BC warm-start halves episodes to reach compliance while preserving final performance, showing offline pre-training accelerates convergence when exploration is costly. Beyond BC, offline RL methods like CQL and IQL mitigate distributional shift, enabling safer policies. A practical approach is a mixed objective: BC for behavior support + conservative RL for improvement without overestimation.
P.3 Practical considerations and compatibility with deployment
Actor–critic and offline pre-training are additive extensions that reuse TempoNet’s slack-quantized embedding representation, permutation-invariant encoder, and lightweight output heads without altering the inference pipeline or latency guarantees. Deployment can follow three modes based on data and compute: a default DQN agent for simplicity, an actor–critic variant for policy-gradient fine-tuning, or a BC-plus-offline-RL pipeline for logged data and limited exploration.
Recommended recipes: with abundant safe logs, apply BC pre-training, offline RL (e.g., CQL or IQL), then short online fine-tuning; for continual adaptation with modest sample complexity, use actor–critic with shared encoder and PPO-style updates, leveraging per-token critic as advantage baseline. Keep actor heads shallow for inference speed and apply standard regularization (weight decay, entropy bonus) for stability. These extensions are orthogonal to TempoNet’s core contributions and can be enabled or disabled without affecting runtime characteristics.
Appendix Q Failure-mode analysis under extreme load
We evaluate TempoNet under an adversarial stress scenario designed to probe the system’s performance boundary. The test parameters are as follows: 600 simultaneously active tasks, mean utilization set to 1.25 with instantaneous peaks up to 1.40, and periodic bursts that inject short high-priority tasks. The injection pattern is ten short tasks every 20 ms over a 10 s interval; each injected task has deadline 10 ms. These conditions produce sustained overload and frequent contention for cores.
Table 19 summarizes aggregate metrics measured across repeated runs (mean $\pm$ std).
Table 19: Adversarial stress-test results. Compliance denotes fraction of tasks meeting deadline; 95th lateness is the 95th percentile of lateness in milliseconds; memory usage is peak resident memory in megabytes.
Method Compliance (%) 95th lateness (ms) Peak memory (MB)
TempoNet $71.3\pm 2.1$ $28.7\pm 3.4$ $312\pm 8$
GNN-based RL[26] $68.9\pm 2.7$ $31.2\pm 4.1$ $295\pm 10$
EDF[27] $52.1\pm 3.0$ $45.6\pm 5.2$
SRPT[25] $63.5\pm 2.5$ $35.1\pm 3.9$
Key observations. TempoNet retains the best overall compliance among compared methods but exhibits an approximately 18 percentage point drop relative to nominal-load performance (for example util = 0.9 settings reported in the main text). Analysis of attention diagnostics reveals a systematic drift under burst conditions: Top-1 alignment falls from 0.92 to 0.79 and average per-step entropy increases from 0.14 to 0.31. These changes indicate that the model concentrates attention more narrowly on extreme low-slack tokens during bursts, which in turn reduces opportunities to schedule longer-deadline tasks and increases deadline miss rates for that cohort.
[…]
To formalize the detection trigger used by runtime mitigations, we monitor the empirical fraction $p$ of tokens whose slack is below the quantization resolution $\Delta$. We compute
$p=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\{s_{i}<\Delta\},$ (100)
where $N$ is the current number of tokens, $s_{i}$ is the slack of token $i$, and $\mathbf{1}\{\cdot\}$ is the indicator function. In this test we observed that when $p>0.30$ the Top-k kernel load increases markedly and attention statistics indicate over-concentration.
Appendix R Runtime mitigations and validated remedies
We implemented two lightweight, online heuristics to limit the observed degradation. Both are purely runtime adaptations that do not require retraining and preserve the core TempoNet design.
Dynamic sparsity scaling. When the system detects the condition $p>\tau_{p}$ with threshold $\tau_{p}=0.30$, it reduces the sparsity budget by scaling the per-block selection parameter $k$. Concretely, the rule sets
$k^{\prime}\;=\;\bigl\lfloor\alpha B\bigr\rfloor,$ (101)
where $\alpha$ is the fraction of block size used for selection and $B$ is block size. In the baseline experiments we use $\alpha_{\text{nom}}=0.10$ and, under the overload trigger, temporarily switch to $\alpha_{\text{burst}}=0.05$. After the burst subsides (measured by $p$ falling below $\tau_{p}$), the system reverts to $\alpha_{\text{nom}}$. This adaptive reduction lowers the effective $k/B$ ratio and prevents the sparse kernel from degrading toward a dense regime.
Dedicated long-slack reserve. One quantization bin is reserved as an insurance bucket for long-slack tasks. Tokens in this bin receive a small positive bias in the encoder, preventing starvation during bursts of ultra-short tasks and preserving execution windows without affecting normal operation.
Empirical impact of mitigations. Applying both mitigations in stress tests restores most lost performance: compliance improves from $71.3\%$ to $76.8\%$, tick overruns drop from $\approx 2\%$ to $0.1\%$, and top-$k$ kernel load returns to nominal. Alignment rises from $0.79$ to $0.86$, entropy falls from $0.31$ to $0.20$, and memory remains stable, indicating no leak or state inflation.
Quantitative summary of the mitigation effect. Let $\Delta_{\mathrm{comp}}$ denote the change in compliance produced by mitigation. In our runs
$\Delta_{\mathrm{comp}}\approx 76.8\%-71.3\%=5.5\%,$ (102)
which bounds the residual performance gap from nominal-load behaviour to within approximately 5 percentage points after applying inexpensive, online heuristics.
Discussion. Stress tests show a clear envelope: when utilization exceeds $\approx 1.2$ and short-task proportion $p>40\%$, TempoNet’s compliance degrades, though not catastrophically. Two lightweight online adaptations, dynamic sparsity scaling and long-slack reserve, mitigate this effectively without retraining, preserving latency guarantees. Recommendations. For deployments facing such bursts, enable both adaptations with conservative thresholds (e.g., $\tau_{p}=0.30$) and hysteresis to prevent oscillation. Persistent overload beyond these regimes requires system-level remedies such as capacity upgrades or admission control.
Appendix S Global Policy Characterization: What TempoNet Learns
This appendix gives a global, human-interpretable summary of the policy learned by TempoNet. It complements attention-based local explanations by deriving a concise rule that approximates the agent’s choices, visualizing state–action preferences, and quantifying how the learned policy differs from common analytic schedulers.
Distilled scheduling rule
We approximate TempoNet’s masked-greedy selections on 600-task industrial traces with a single-line deterministic priority rule. At each decision step tasks are ordered by priority $\text{priority}=\alpha\cdot\big(1/(\tilde{s}+1)\big)+\beta\cdot\big(1/(c+1)\big)$, where $\tilde{s}$ denotes the quantized slack index and $c$ the remaining execution. A grid search over $(\alpha,\beta)$ returns normalized weights $\alpha=0.73$ and $\beta=0.27$. Applying this linear rule reproduces the agent’s masked-greedy choices on 91% of decisions, indicating that, at the global level, TempoNet behaves like a weighted combination of minimum-slack and shortest-remaining-processing-time priorities.
Figure 12: Probability of selection conditioned on quantized slack and normalized remaining execution.
State–action preference heat map
Figure 12 reports the empirical selection probability as a function of quantized slack (horizontal axis) and normalized remaining execution (vertical axis), aggregated over roughly 600k decisions using 20 ms sampling bins. The densest (darkest) region lies in the lower-left quadrant, showing that tasks with both small slack and small remaining work are chosen most often, which provides evidence that the policy jointly accounts for urgency and residual cost.
Policy distance to EDF and SRPT
We compare action sequences produced on the same trace by TempoNet, a pure EDF scheduler (deadline-only), and a pure SRPT scheduler (remaining-time-only). Agreement is measured as the fraction of identical actions (Hamming agreement) between two sequences. Results are summarized in Table 20.
Table 20: Action-sequence agreement between TempoNet and classical schedulers or the distilled linear rule.
Pair Action agreement Interpretation
TempoNet vs. EDF 68% leans toward shorter jobs relative to EDF
TempoNet vs. SRPT 71% gives additional weight to deadlines vs. SRPT
TempoNet vs. distilled rule 91% closely matched by a single weighted rule
Taken together, these results show that although TempoNet is learned via reinforcement learning, its emergent global policy is well approximated by a transparent, weighted slack rule that balances deadline urgency and remaining execution. This characterization addresses interpretability concerns and helps explain why TempoNet often outperforms pure EDF or SRPT on deadline-focused metrics.

### v2

@@
 Earliest Deadline First (EDF) 11.67%
-Feedforward DQN (FF-DQN) 71.43%
+Feedforward DQN ($F$-DQN) 71.43%
 TempoNet (Proposed) 79.00% +7.57%
@@
 We conduct comprehensive evaluations across three computational scenarios: uniprocessor periodic scheduling with standardized task configurations, industrial multi-core workloads with mixed-criticality workflows, and large-scale multiprocessor systems with 100–600 tasks.
-Our framework is benchmarked against established scheduling approaches including classical schedulers (RM, EDF), feedforward Deep Q-Network (FF-DQN), Dynamic Importance-aware Online Scheduling (DIOS), and quantum-inspired optimization methods.
+Our framework is benchmarked against established scheduling approaches including classical schedulers (RM, EDF), feedforward Deep Q-Network ($F$-DQN), Dynamic Importance-aware Online Scheduling (DIOS), and quantum-inspired optimization methods.
 Evaluation Metrics: Performance assessment employs deadline compliance rate, average response time, and computational overhead.
@@
 A.1 Definitions and Policy Formalization
-Let $\mathbf{s}=[s_{1},\dots,s_{N}]^{\top}$ denote a vector representing the temporal laxity (slack) of $N$ active tasks, where each coordinate $s_{i}$ is bounded within the compact interval $0,S_{\max}]$.
+Let $\mathbf{s}=[s_{1},\dots,s_{N}]^{\top}$ denote a vector representing the temporal laxity (slack) of NN active tasks, where each coordinate $s_{i}$ is bounded within the compact interval $0,S_{\max}]$.
 We consider two distinct families of scheduling policies:
@@
 $\Theta\bigl(N\log N+m\bigr),$ (45)
-where $m$ is the number of cores to fill and $N$ is the number of available tasks.
+where $m$ is the number of cores to fill and NN is the number of available tasks.
 where the cost expressions above quantify the primary algorithmic operations: an argsort over the score vector and $m$ trivial mask updates.
@@
 Constructive counterexample (SRPT can miss deadlines that a slack-based policy satisfies).
-Consider two tasks that arrive at time $t=0$ with the following parameters: task $A$ has remaining processing $c_{A}=1$ and deadline $d_{A}=100$; task $B$ has remaining processing $c_{B}=2$ and deadline $d_{B}=2.2$.
-SRPT schedules the shorter job $A$ first, finishing it by $t=1$, then executes $B$ and completes $B$ at $t=3$, which misses $B$’s deadline.
+Consider two tasks that arrive at time $t=0$ with the following parameters: task AA has remaining processing $c_{A}=1$ and deadline $d_{A}=100$; task $B$ has remaining processing $c_{B}=2$ and deadline $d_{B}=2.2$.
+SRPT schedules the shorter job AA first, finishing it by $t=1$, then executes $B$ and completes $B$ at $t=3$, which misses $B$’s deadline.
 A slack-minimizing policy computes initial slacks $s_{A}=99$ and $s_{B}=0.2$ and therefore schedules $B$ first, completing both tasks before their deadlines.
@@
 Q4 (longest) $180\pm 35$ $85.9\pm 2.7$
-KS-test $p$-value 0.18 (no bias)
+KS-test pp-value 0.18 (no bias)
 Table 9 reports a small variation in meet rates across quartiles (the range is approximately 1.3 percentage points).
@@
 A: masked-greedy 90.1 12.4 420 default, sub-ms
-B: Sinkhorn 90.6 12.1 860 +0.5 pp, $\times$2 latency
+B: Sinkhorn 90.6 12.1 860 +0.5 $p$, $\times$2 latency
 K.2 Shaped-reward Ablation
@@
 600 298 $\pm$ 12 71 $\pm$ 5 375 $\pm$ 17 $\sim$430
-Figure 9: Real-time boundary: synthetic inference latency versus number of active tasks $N$ with a 1 ms budget line.
+Figure 9: Real-time boundary: synthetic inference latency versus number of active tasks NN with a 1 ms budget line.
 Vertical dashed line marks the interpolated upper bound $N_{\max}$.
@@
 The attention scores are sparsified by retaining only the top-k values per query within predefined blocks.
-Let the attention score matrix be $A\in\mathbb{R}^{(N+1)\times(N+1)}$, where $N$ is the number of tasks.
+Let the attention score matrix be $A\in\mathbb{R}^{(N+1)\times(N+1)}$, where NN is the number of tasks.
 The matrix is partitioned into blocks of size $B\times B$, where $B$ is the block size determined based on the task count and hardware constraints.
@@
 The theoretical complexity of the sparse attention mechanism is $O(N^{1.1})$ on average, achieved through the combination of block Top-k and chunking.
-To validate this, we measured the wall-clock time for attention computation across task sets of size $N$ ranging from 10 to 600 tasks.
+To validate this, we measured the wall-clock time for attention computation across task sets of size NN ranging from 10 to 600 tasks.
 The results, plotted in Figure 1, show that the time $T(N)$ fits the model $T(N)=c\cdot N^{1.1}+d$, where $c$ and $d$ are constants determined via linear regression on log-transformed data.
@@
 $s_{t}=\left(\tilde{s}_{1}(t),\tilde{s}_{2}(t),\ldots,\tilde{s}_{N}(t),a_{c}(t)\right)$ (81)
-where $\tilde{s}_{i}(t)$ refers to the quantized slack index of task $i$ at time $t$ as per Equation (5) in the main text, representing the task’s urgency level, and $a_{c}(t)\in\{1,2,\ldots,m\}^{N}$ is a core assignment vector where each component $a_{c}^{(i)}(t)$ indicates the core index to which task $i$ is currently assigned, with $m$ being the number of cores and $N$ the total number of tasks.
+where $\tilde{s}_{i}(t)$ refers to the quantized slack index of task $i$ at time $t$ as per Equation (5) in the main text, representing the task’s urgency level, and $a_{c}(t)\in\{1,2,\ldots,m\}^{N}$ is a core assignment vector where each component $a_{c}^{(i)}(t)$ indicates the core index to which task $i$ is currently assigned, with $m$ being the number of cores and NN the total number of tasks.
 L.5.2 Action Space Definition
@@
 Notation and setting.
-Consider one self-attention layer applied to a sequence of $N$ tokens.
+Consider one self-attention layer applied to a sequence of NN tokens.
 The sequence is partitioned into $m$ non-overlapping blocks of equal size $B$, so that $N=mB$.
@@
 where $C(N;B,k)$ is the total number of query–key score evaluations per head per layer, $m=N/B$ is the number of blocks, and $\mathrm{cost}_{\mathrm{select}}(B)$ denotes the cost to compute block summaries and choose top-$k$ candidate blocks for a single block.
-The three terms on the right-hand side correspond respectively to intra-block pairwise scores, the aggregated cost of block selection across all blocks, and inter-block score evaluations incurred when each query inspects up to $kB$ external keys.
+The three terms on the right-hand side correspond respectively to intra-block pairwise scores, the aggregated cost of block selection across all blocks, and inter-block score evaluations incurred when each query inspects up to kBkB external keys.
 Simplified estimate under light-weight selection.
@@
 Practical parameter regimes and their implications.
-If both $B$ and $k$ are constants independent of $N$, the complexity in (86) is linear in $N$.
-If $B$ and $k$ scale like $\log N$, the cost grows polylogarithmically times $N$.
+If both $B$ and $k$ are constants independent of NN, the complexity in (86) is linear in NN.
+If $B$ and $k$ scale like $\log N$, the cost grows polylogarithmically times NN.
 If $B=\Theta(\sqrt{N})$ the dominant contributions scale as $N^{3/2}$ and the per-layer cost becomes super-linear.
@@
 where $\mathbb{E}[\#\mathrm{nz}]$ stands for the expected total number of retained (non-zero) attention entries per layer under the assumed input distribution and selection policy.
-Substituting $C=\Theta(N/\log N)$, $B=\Theta(\sqrt{N})$ and $k=\Theta(\sqrt{N})$ and simplifying shows that the intra-chunk term dominates for sufficiently large $N$, yielding the asymptotic estimate
+Substituting $C=\Theta(N/\log N)$, $B=\Theta(\sqrt{N})$ and $k=\Theta(\sqrt{N})$ and simplifying shows that the intra-chunk term dominates for sufficiently large NN, yielding the asymptotic estimate
 $\mathbb{E}[\#\mathrm{nz}]\;=\;\Theta\!\Bigl(\frac{N^{1.5}}{\sqrt{\log N}}\Bigr),$ (92)
@@
 First, symmetric sparsity patterns inside blocks allow optimized GPU kernels to fuse row- and column-wise accesses, reducing effective memory traffic and lowering constant factors.
-Second, fitting a power law to wall-clock times over a finite $N$ range, combined with caching and kernel fusion, often produces an apparent exponent smaller than the asymptotic one.
+Second, fitting a power law to wall-clock times over a finite NN range, combined with caching and kernel fusion, often produces an apparent exponent smaller than the asymptotic one.
 Empirically, these practicalities can transform the theoretical $N^{1.5}$ scaling into a measured behaviour close to $N^{1+\varepsilon}$ with small $\varepsilon$; the experiments reported in the paper fitted an exponent near $1.1$ with high goodness-of-fit.
@@
 $C(N)=\underbrace{NB}_{\text{intra-block}}\;+\;\underbrace{m\cdot\mathrm{cost}_{\mathrm{select}}(B)}_{\text{selection overhead}}\;+\;\underbrace{NkB}_{\text{inter-block}},$ (93)
-where the meaning of each symbol is as stated above. Fixing $B$ and $k$ keeps $C(N)$ linear in $N$; allowing either to grow with $N$ can push the cost into super-linear regimes.
+where the meaning of each symbol is as stated above. Fixing $B$ and $k$ keeps $C(N)$ linear in NN; allowing either to grow with NN can push the cost into super-linear regimes.
 Table 17: Per-layer attention complexity under different regimes.
@@
 Block Top-k with $B=\Theta(\sqrt{N}),k=\Theta(\sqrt{N})$ $\Theta\!\bigl(N^{1.5}/\sqrt{\log N}\bigr)$. Average-case analytic estimate under chunking.
-Measured (optimized CUDA kernel, finite $N$ range) $\Theta(N^{1.1})$ (empirical fit). Observed when kernel fusion and symmetric sparsity reduce constants.
+Measured (optimized CUDA kernel, finite NN range) $\Theta(N^{1.1})$ (empirical fit). Observed when kernel fusion and symmetric sparsity reduce constants.
 Summary
@@
 $r(t)=\sum_{i=1}^{N}\left(I\{c_{i}(t-1)>0\wedge c_{i}(t)=0\}-I\{t=d_{i}^{(k)}\wedge c_{i}(t)>0\}\right)$ (94)
-where $I\{\cdot\}$ denotes the indicator function that equals 1 if the enclosed condition is true and 0 otherwise, $c_{i}(t)$ represents the remaining execution time of task $i$ at time $t$, and $d_{i}^{(k)}$ is the absolute deadline of the $k$-th job instance of task $i$.
+where $I\{\cdot\}$ denotes the indicator function that equals 1 if the enclosed condition is true and 0 otherwise, $c_{i}(t)$ represents the remaining execution time of task $i$ at time $t$, and $d_{i}^{(k)}$ is the absolute deadline of the $k$-th job instance of task ii.
 The first term rewards the completion of a job within the current time step, while the second term penalizes a deadline miss for any active job.
@@
 $p=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\{s_{i}<\Delta\},$ (100)
-where $N$ is the current number of tokens, $s_{i}$ is the slack of token $i$, and $\mathbf{1}\{\cdot\}$ is the indicator function.
+where NN is the current number of tokens, $s_{i}$ is the slack of token $i$, and $\mathbf{1}\{\cdot\}$ is the indicator function.
 In this test we observed that when $p>0.30$ the Top-k kernel load increases markedly and attention statistics indicate over-concentration.